### v1

# An $H^{1}$ multiplier theorem on anisotropic Hardy sapces

Abstract.
A parallel result of Sledd–Stegenga’s $H^{1}\rightarrow L^{1}$ multiplier theorem was obtained on the $H^{1}$ space under the anisotropic settings. Based on the same technique, an $H^{1}\rightarrow L^{p}$ multiplier theorem is also proved for $1\leq p<\infty$.
1. Introduction
[…]
$\Big{(}\int_{\mathbb{R}^{d}}\frac{|\widehat{f}|^{p}}{|\xi|^{(2-p)d}}\,\mathrm{d}\xi\Big{)}^{\frac{1}{p}}\lesssim_{p,d}\|f\|_{H^{p}(\mathbb{R}^{d})}.$
When $p=1$, we have a stronger result, due to Sledd and Stegenga (see [8], Theorem 1).
The $H^{1}$ boundedness criterion.
Let $\mu$ be a positive Borel measure on $\mathbb{R}^{d}\setminus\{0\}$, then
$\int_{\mathbb{R}^{d}}|\widehat{f}(\xi)|\,\mathrm{d}\mu(\xi)\lesssim\|f\|_{H^{1}(\mathbb{R}^{d})},$
if and only if
$\sup_{\epsilon>0}\Big{(}\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{0\}}\mu\big{(}Q_{\epsilon}(\alpha)\big{)}^{2}\Big{)}^{\frac{1}{2}}<\infty,$
where $Q_{\epsilon}(\alpha)$ is the cube $\{(x_{i})_{i=1}^{d}\in\mathbb{R}^{d}:\epsilon\alpha_{i}-\frac{\epsilon}{2}\leq x_{i}<\epsilon\alpha_{i}+\frac{\epsilon}{2}\}$.
It is known that, under the anisotropic settings $H^{p}_{A}$ (we assume $0<p\leq 1$), the Hardy–Littlewood inequality still holds true: Let $A$ be a dilation matrix, and $\rho_{\ast}(\cdot)$ is a quasi-norm associated with $A^{\ast}$, then
$\Big{(}\int_{\mathbb{R}^{d}}\frac{|\widehat{f}|^{p}}{\rho_{\ast}(\xi)^{2-p}}\,\mathrm{d}\xi\Big{)}^{\frac{1}{p}}\lesssim_{p,d,A}\|f\|_{H^{p}_{A}(\mathbb{R}^{d})}.$
This inequality was proved by Bownik–Wang, see [3], Corollary 8.
However, the anisotropic analog of Sledd–Stegenga’s criterion seems to be absent so far. We will close this gap. The statement of the criterion under the anisotropic settings is more complicated and requires more preparation. Moreover, some arguments in [8] does not work any more, and we need few modifications.
Also, there are problems in the original proofs of [8], which either omit many non-trivial steps or are too specific (only the 1-dimensional case is proved). Here we complete the missing details of the proof in [8].
2. Preliminaries
In this section, most of the text is adapted from the book [2].
2.1. Dilation matrix and its homogeneous norm
Let $A\in\mathbf{GL}(d,\mathbb{R})$ and $b\coloneqq|\det(A)|$. We say that $A$ is a dilation matrix, if all its eigenvalues $\{\lambda_{i}\}_{i=1}^{d}$ have modulus strictly greater than $1$.
For each dilation matrix $A$, there exists a measurable function (not unique) $\rho_{A}\colon\mathbb{R}^{d}\rightarrow[0,\infty)$, called a homogenous quasi-norm associated with $A$, satisfying the following properties:
  1. (1)
    $\rho_{A}(x)=0$ if and only if $x=0$.
  2. (2)
    $\rho_{A}(Ax)=b\rho_{A}(x)$ for all $x\in\mathbb{R}^{d}$.
  3. (3)
    $\rho_{A}(x+y)\leq c(\rho_{A}(x)+\rho_{A}(y))$ for all $x,y\in\mathbb{R}^{d}$, where $c>0$ is an absolute constant independent of $x$ and $y$.
We choose $\lambda_{-}$ and $\lambda_{+}$ so that $1<\lambda_{-}<|\lambda_{1}|\leq\ldots\leq|\lambda_{d}|<\lambda_{+}$. The following growth estimate of $\rho_{A}$ is known.
Proposition 2.1.
(see [6], Appendix B) Suppose that $\rho_{A}$ is a homogenous quasi-norm associated with the matrix $A$, then there is a constant $c_{A}>0$, so that
$c_{A}^{-1}\rho_{A}(x)^{\zeta_{-}}$ $\leq|x|\leq c_{A}\rho_{A}(x)^{\zeta_{+}},\text{ if }\rho_{A}(x)\geq 1,$
$c_{A}^{-1}\rho_{A}(x)^{\zeta_{+}}$ $\leq|x|\leq c_{A}\rho_{A}(x)^{\zeta_{-}},\text{ if }\rho_{A}(x)<1.$
Here $\zeta_{\pm}\coloneqq\ln\lambda_{\pm}/\ln b$, and $c_{A}$ only depends on $\zeta_{\pm}$.
The basic geometric objects in isotropic settings are Euclidean balls $B(x,r)$. For the anisotropic case, we have the following substitute. We say that $\Delta\subset\mathbb{R}^{d}$ is an ellipse, if there exists $Q\in\mathbf{GL}(d,\mathbb{R})$, so that $\Delta=Q^{-1}(B(0,1))$, where $B(0,1)$ is the (open) unit Euclidean ball in $\mathbb{R}^{d}$.
Proposition 2.2.
(See [2], Chapter 1, Lemma 2.2) Suppose that $A$ is a dilation matrix, then there exists an ellipse $\Delta\subset\mathbb{R}^{d}$, so that
$\Delta\subset A(\Delta).$
This proposition is similar to the fact that $B(0,r)\subset B(0,R)$ whenever $r<R$. By rescaling, we may assume that $|\Delta|\approx 1$. Although $\Delta$ is not unique, we will choose a specific one and refer it as the ellipse associated with $A$.
The family of balls $\{\Delta_{k}\}_{k\in\mathbb{Z}}$ associated with $A$ around the origin is defined by
$\Delta_{k}\coloneqq A^{k}(\Delta),\text{ where }k\in\mathbb{Z}.$
By Proposition 2.2, we have $\Delta_{k}\subset\Delta_{k+1}$, and $|\Delta_{k}|=b^{k}|\Delta_{0}|\approx b^{k}$.
Although all quasi-norms $\rho_{A}$ are equivalent (see [2], Chapter 1, Lemma 2.4), we will use the following specific one: From the nested property of the family $\{\Delta_{k}\}_{k\in\mathbb{Z}}$, we know that for every $x\in\mathbb{R}^{d}\setminus\{0\}$, there exists a unique $k\in\mathbb{Z}$ so that $x\in\Delta_{k+1}\setminus\Delta_{k}$. We define
$\rho_{A}(x)\coloneqq\begin{cases}b^{k}&\quad\text{if }x\in\Delta_{k+1}\setminus\Delta_{k},\\ 0&\quad\text{otherwise}.\end{cases}$
Let $A^{\ast}$ be the adjoint of $A$, since $A^{\ast}$ and $A$ have the same spectrum, the matrix $A^{\ast}$ is also a dilation matrix. We can define the quasi-norm $\rho_{A^{\ast}}$, the nested balls $\{\Delta^{\ast}_{k}\}_{k\in\mathbb{Z}}$, and so on.
2.2. Anisotropic $H^{1}$, and the atomic decomposition
Let $\varphi\in\mathcal{S}(\mathbb{R}^{d})$ with $\int\varphi\neq 0$. The anisotropic dilation of $\varphi$ is defined by $\varphi_{k}(\cdot)\coloneqq b^{k}\varphi(A^{k}\cdot)$. For $f\in L^{1}(\mathbb{R}^{d})$, the smooth maximal function of $f$ with respect to $\varphi$ is defined by
$\mathcal{M}_{\varphi}(f)(x)\coloneqq\sup_{k\in\mathbb{Z}}|f\ast\varphi_{k}|(x).$
The anisotropic $H^{1}$ space associated with the dilation $A$ is defined by
$H^{1}_{A}(\mathbb{R}^{d})\coloneqq\Big{\{}f\in L^{1}(\mathbb{R}^{d}):\mathcal{M}_{\varphi}(f)\in L^{1}(\mathbb{R}^{d})\Big{\}}$
Similar to the isotropic case, the space $H^{1}_{A}$ can be characterized by the atomic decomposition. We say that an $L^{1}$ function $a$ is an $H^{1}_{A}$-atom, if
  1. (1)
    There exist $x_{0}\in\mathbb{R}^{d}$ and $k\in\mathbb{Z}$, so that $\mathrm{supp}(a)\subset x_{0}+\Delta_{k}$.
  2. (2)
    The function $a$ is bounded by $|\Delta_{k}|^{-1}$ almost everywhere: $\|a\|_{\infty}\leq|\Delta_{k}|^{-1}$.
  3. (3)
    The function $a$ has vanishing moment: $\int a=0$.
The atomic decomposition of $H^{1}_{A}$ states that $f\in H^{1}_{A}$ if and only if there exists atoms $\{a_{i}\}_{i=1}^{\infty}$ and a sequence $(\lambda_{i})_{i=1}^{\infty}\in\ell^{1}$, so that $f=\sum_{i=1}^{\infty}\lambda_{i}a_{i}$. Moreover,
$\|f\|_{H^{1}_{A}}\approx\inf\Big{\{}\sum_{i=1}^{\infty}|\lambda_{i}|:f=\sum_{i=1}^{\infty}\lambda_{i}a_{i}\Big{\}},$
where the infimum is taken over all possible atomic decomposition of $f$. For this characterization, see [3], Chapter 1, Theorem 6.5.
2.3. Duality, and anisotropic $\mathrm{BMO}$
We denote $\mathcal{B}$ the family of translated balls associated to $A$,
$\mathcal{B}\coloneqq\{x+\Delta_{k}:x\in\mathbb{R}^{d},\text{ and }k\in\mathbb{Z}\}.$
Assume that $1\leq q<\infty$. A locally integrable function $f$ is said to be in the anisotropic bounded mean oscillation space $\mathrm{BMO}_{A,q}$, if the following inequality holds:
$\sup_{B\in\mathcal{B}}\inf_{c_{B}\in\mathbb{C}}\bigg{(}\frac{1}{|B|}\int_{B}|f(x)-c_{B}|^{q}\,\mathrm{d}x\bigg{)}^{\frac{1}{q}}<\infty.$
This quantity also defines the $\mathrm{BMO}_{A,q}$ norm. It is well-known that $\mathrm{BMO}_{A,q}$ norms are equivalent for $1\leq q<\infty$, so we will use the abbreviation $\mathrm{BMO}_{A}$ and omit $q$. Also, as in the isotropic case, the following duality relation holds:
$(H^{1}_{A})^{\ast}=\mathrm{BMO}_{A}.$
For basic properties of the anisotropic $\mathrm{BMO}$, see [2], Chapter 1, Section 8.
2.4. Use of notations and abbreviations
For the later part of this article, we will drop the matrix $A$ and $A^{\ast}$ in the notations, which does not create ambiguity. For example, we use $\rho,\rho_{\ast}$ to denote $\rho_{A},\rho_{A^{\ast}}$ respectively. Unless otherwise specified, the spaces $H^{1}$ and $\mathrm{BMO}$ are always interpreted as $H^{1}_{A}$ and $\mathrm{BMO}_{A}$.
3. Main theorem and its proof
To state the main theorem, we need to introduce few more notations. For $\Delta\subset\mathbb{R}^{d}$ the ellipse associated with $A$, we define $R=R(\Delta)$ be the smallest rectangle contains $\Delta$. Here, for rectangles, we mean that the sides of $R$ are parallel to the coordinate axes. By simple convex geometry, we know that $|\Delta|\approx|R|$, and there exists $M=M_{\Delta}\in\mathbb{N}$, so that
$A^{-M}(R)\subset\Delta\subset R.$
For $\Delta^{\ast}$, similar conclusion holds. We denote the smallest rectangle containing $\Delta^{\ast}$ by $R^{\ast}$, and assume that $(A^{\ast})^{-N}(R^{\ast})\subset\Delta^{\ast}\subset R^{\ast}$, where $N\in\mathbb{N}$ depends on the ellipse $\Delta^{\ast}$, and the matrix $A^{\ast}$.
Suppose that $R=I_{1}\times I_{2}\cdots\times I_{d}$. For $\alpha\in\mathbb{Z}^{d}$, We denote $R(\alpha)$ be the translation
$R(\alpha)\coloneqq\Big{\{}(x_{i})_{i=1}^{d}:|I_{i}|\alpha_{i}-\frac{|I_{i}|}{2}\leq x_{i}<|I_{i}|\alpha_{i}+\frac{|I_{i}|}{2}\Big{\}},$
and $\Delta(\alpha)\subset R(\alpha)$ be the corresponding translation of $\Delta$.
Theorem 3.1.
Let $\mu$ be a positive Borel measure on $\mathbb{R}^{d}\setminus\{0\}$. Then
(3.1) $\int_{\mathbb{R}^{d}}|\widehat{f}|\,\mathrm{d}\mu\lesssim\|f\|_{H^{1}},$
if and only if
(3.2) $\sup_{k\in\mathbb{Z}}\left(\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{0\}}\mu\left(\left(A^{\ast}\right)^{k}\left(R^{\ast}\left(\alpha\right)\right)^{2}\right)\right)^{\frac{1}{2}}\lesssim 1.$
3.1. Proof of the sufficient part
Assume that (3.2) holds. It suffices to prove the result for $H^{1}$-atoms $a$. Furthermore, since we consider the magnitude of the Fourier transform, by the identity $\mathcal{F}(a(x-x_{0}))(\xi)=\mathrm{e}^{-2\pi\mathrm{i}x_{0}\cdot\xi}\widehat{a}(\xi)$, we may assume that $a$ is supported in $\Delta_{k}$. We decompose $\int|\widehat{a}|\,\mathrm{d}\mu$ into two parts: the integral over the region $(A^{\ast})^{-(k+N)}(R^{\ast})$, and over its complement.
Let us begin with the integral on $(A^{\ast})^{-(k+N)}(R^{\ast})$. It is easily seen that
$\int_{(A^{\ast})^{-(k+N)}(R^{\ast})}|\widehat{a}(\xi)|\,\mathrm{d}\mu\lesssim\sum_{l=-\infty}^{-(k+N)}\int_{(A^{\ast})^{l}(R^{\ast})\setminus(A^{\ast})^{l-N}(R^{\ast})}|\widehat{a}(\xi)|\,\mathrm{d}\mu.$
As a consequence, to control $\int_{(A^{\ast})^{-(k+N)}(R^{\ast})}|\widehat{a}(\xi)|\,\mathrm{d}\mu$, we only need to obtain good estimates of $\int_{(A^{\ast})^{l}(R^{\ast})\setminus(A^{\ast})^{l-N}(R^{\ast})}|\widehat{a}|\,\mathrm{d}\mu$, for each $l\leq-(k+N)$.
The following finite covering property is straightforward to verify.
Proposition 3.2.
The region $(A^{\ast})^{l}(R^{\ast})\setminus(A^{\ast})^{l-N}(R^{\ast})$ can be covered by $\mathcal{O}_{N}(1)$ translations of $(A^{\ast})^{l-N}(R^{\ast}(\alpha))$ with $\alpha\neq(0,0)$.
This figure explains the proposition for $R^{\ast}\setminus(A^{\ast})^{-N}(R^{\ast})$. The rectangle drawn with bold lines is $R^{\ast}$. The small parallelogram centered at the origin is $(A^{\ast})^{-N}(R^{\ast})$, and 26 translations of it cover the region $R^{\ast}\setminus(A^{\ast})^{-N}(R^{\ast})$. The general case follows immediately by applying the mapping $(A^{\ast})^{l}$.
Figure 1.
Using Proposition 3.2 and the Cauchy–Schwarz, it follows that
$\mu\Big{(}(A^{\ast})^{l}(R^{\ast})\setminus(A^{\ast})^{l-N}(R^{\ast})\Big{)}$ $\leq\sum_{\alpha:\,\mathcal{O}_{N}(1)\textrm{ terms}}\mu\Big{(}(A^{\ast})^{l-N}(R^{\ast}(\alpha))\Big{)}$
$\lesssim\mathcal{O}_{N}(1)\bigg{(}\sum_{\alpha:\,\mathcal{O}_{N}(1)\textrm{ terms}}\mu\bigg{(}(A^{\ast})^{l-N}(R^{\ast}(\alpha))\bigg{)}^{2}\bigg{)}^{\frac{1}{2}}$
$\lesssim\mathcal{O}_{N}(1)\bigg{(}\sum_{\alpha\neq 0}\mu\bigg{(}(A^{\ast})^{l-N}(R^{\ast}(\alpha))\bigg{)}^{2}\bigg{)}^{\frac{1}{2}}$
$\lesssim_{N}1,$
where we used the hypothesis (3.2) in the last line.
We recall that $\Delta^{\ast}\subset R^{\ast}\subset(A^{\ast})^{N}(\Delta^{\ast})$. Applying the mapping $(A^{\ast})^{l-N}$ leads to
$(A^{\ast})^{l-N}(R^{\ast})\subset(A^{\ast})^{l}(\Delta^{\ast})\subset(A^{\ast})^{l}(R^{\ast})\subset(A^{\ast})^{l+N}(\Delta^{\ast}),$
and it follows trivially that $(A^{\ast})^{l}(R^{\ast})\setminus(A^{\ast})^{l-N}(R^{\ast})\subset(A^{\ast})^{l+N}(\Delta^{\ast})$.
The analysis above shows that, if $l\leq-(k+N)$, then
$\rho_{\ast}(\xi)\leq b^{l+N}\leq b^{-k},\text{ whenever }\xi\in(A^{\ast})^{l}(R^{\ast})\setminus(A^{\ast})^{l-N}(R^{\ast}).$
We will use the following Fourier transform estimate of $H^{1}$-atoms, which was established by Bownik–Wang:
Proposition 3.3.
(See [3], Lemma 5) Suppose that $a$ is an $H^{1}$-atom supported in $\Delta_{k}$, then
$|\widehat{a}(\xi)|\lesssim b^{k\zeta_{-}}\rho_{\ast}(\xi)^{\zeta_{-}},\text{ whenever }\rho_{\ast}(\xi)\leq b^{-k}.$
Combining previous estimates, we can write
$1$ $\gtrsim\mu\Big{(}(A^{\ast})^{l}(R^{\ast})\setminus(A^{\ast})^{l-N}(R^{\ast})\Big{)}$
$=\int_{(A^{\ast})^{l}(R^{\ast})\setminus(A^{\ast})^{l-N}(R^{\ast})}\frac{1}{\rho_{\ast}^{\zeta_{-}}(\xi)}\rho_{\ast}^{\zeta_{-}}(\xi)\,\mathrm{d}\mu$
$\gtrsim\frac{1}{b^{(l+N)\zeta_{-}}}\int_{(A^{\ast})^{l}(R^{\ast})\setminus(A^{\ast})^{l-N}(R^{\ast})}b^{-k\zeta_{-}}|\widehat{a}(\xi)|\,\mathrm{d}\mu$
$=\frac{1}{b^{(l+k+N)\zeta_{-}}}\int_{(A^{\ast})^{l}(R^{\ast})\setminus(A^{\ast})^{l-N}(R^{\ast})}|\widehat{a}(\xi)|\,\mathrm{d}\mu.$
Summing over $l\leq-(k+N)$ shows that
$\int_{(A^{\ast})^{-(k+N)}(R^{\ast})}|\widehat{a}(\xi)|\,\mathrm{d}\mu$ $\lesssim\sum_{l=-\infty}^{-(k+N)}\int_{(A^{\ast})^{l}(R^{\ast})\setminus(A^{\ast})^{l-N}(R^{\ast})}|\widehat{a}(\xi)|\,\mathrm{d}\mu$
$\lesssim\sum_{l=-\infty}^{-(k+N)}b^{(l+k+N)\zeta_{-}}$
$\lesssim 1.$
The implicit constant depends (of course) on $N,b$ and $\zeta_{-}$, but not on $k$.
Remark 1.
In the original paper [8], the estimate over $(A^{\ast})^{-(k+N)}(R^{\ast})$ is omitted and the authors said it was an easy consequence of (3.2). This is not quite accurate. In [8], the authors write the summation (3.2) over all $\alpha\in\mathbb{Z}^{d}$, which trivially implies the estimate over $(A^{\ast})^{-(k+N)}(R^{\ast})$. However, the correct condition is the sum over $\alpha\in\mathbb{Z}^{d}\setminus\{0\}$. In this case, we need few more work.
Having completed the estimate of the integral on $(A^{\ast})^{-(k+N)}(R^{\ast})$, we now turn to control the integral on the complement $\mathbb{R}^{d}\setminus(A^{\ast})^{-(k+N)}(R^{\ast})$, it is easy to check that $\mathbb{R}^{d}\setminus R^{\ast}=\bigcup_{\alpha\in\mathbb{Z}^{d}\setminus\{(0,0)\}}R^{\ast}(\alpha)$, applying the invertible mapping $(A^{\ast})^{-(k+N)}$ gives
$\mathbb{R}^{d}\setminus(A^{\ast})^{-(k+N)}(R^{\ast})=\bigcup_{\alpha\in\mathbb{Z}^{d}\setminus\{(0,0)\}}(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha)),$
and the union is a disjoint union. Therefore
$\int_{\mathbb{R}^{d}\setminus(A^{\ast})^{-(k+N)}(R^{\ast})}|\widehat{a}|\,\mathrm{d}\mu$ $=\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{(0,0)\}}\int_{(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))}|\widehat{a}|\,\mathrm{d}\mu$
[…]
By our hypothesis (3.2),
$\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{(0,0)\}}\mu\Big{(}(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))\Big{)}^{2}\lesssim 1.$
By Cauchy–Schwarz, the proof will be completed if we can show that
(3.3) $\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{(0,0)\}}\sup_{(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))}|\widehat{a}|^{2}\lesssim 1.$
After a change of variable $\eta=(A^{\ast})^{k+N}\xi$, we have
$\sup_{\xi\in(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))}|\widehat{a}(\xi)|^{2}=\sup_{\eta\in R^{\ast}(\alpha)}\Big{|}\widehat{a}\Big{(}(A^{\ast})^{-(k+N)}\eta\Big{)}\Big{|}^{2}.$
The basic properties of the Fourier transform implies that
$\widehat{a}\Big{(}(A^{\ast})^{-(k+N)}\eta\Big{)}=b^{k+N}\mathcal{F}\big{(}a\circ A^{k+N}\big{)}(\eta).$
We are reduced to control the sum
$\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{(0,0)\}}\sup_{(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))}|\widehat{a}|^{2}=b^{2(k+N)}\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{(0,0)\}}\sup_{\eta\in R^{\ast}(\alpha)}\Big{|}\mathcal{F}\big{(}a\circ A^{k+N}\big{)}(\eta)\Big{|}^{2}$
Proposition 3.4.
Assume that $d=2$. Suppose that $f$ is smooth, and $Q=I_{1}\times I_{2}$ is a rectangle in $\mathbb{R}^{2}$. The following inequality holds:
$\sum_{\alpha\in\mathbb{Z}^{2}}\sup_{Q(\alpha)}|f|^{2}\lesssim\frac{|I_{1}|}{|I_{2}|}\int|\partial_{1}f|^{2}+|Q|\int|\partial_{2}\partial_{1}f|^{2}+\frac{|I_{2}|}{|I_{1}|}\int|\partial_{2}f|^{2}+\frac{1}{|Q|}\int|f|^{2},$
where the implicit constant in $\lesssim$ is independent of $Q$.
Remark 2.
This result was proved by Sledd-Stegenga for $d=1$ (see [8], Theorem 3), which states that for any finite interval $I\subset\mathbb{R}$,
$\sum_{\alpha\in\mathbb{Z}}\sup_{I(\alpha)}|f|^{2}\leq 2\bigg{(}\frac{1}{|I|}\int_{\mathbb{R}}|f|^{2}+|I|\int_{\mathbb{R}}|f^{\prime}|^{2}\bigg{)}.$
In [8], the authors wrote that the higher-dimensional case is somewhat complicated so the proof was omitted. We give a proof of the two-dimensional case. For general $d\geq 3$, the proof below also works, but the notations will become more annoyed.
proof of Proposition 3.4.
Write $b_{Q}\coloneqq\frac{1}{|Q|}\int_{Q}f(x)\ \mathrm{d}x$, then $b_{Q}$ is independent of $x$, and
$\sup_{Q}|f|^{2}\leq 2\sup_{Q}|f-b_{Q}|^{2}+2|b_{Q}|^{2}.$
Let us first see how to control $|b_{Q}|^{2}$, which is easier. The Cauchy–Schwarz inequality implies $|b_{Q}|^{2}\leq\frac{1}{|Q|}\int_{Q}|f|^{2}$. Summing over all $\alpha\in\mathbb{Z}^{2}$, we obtain
$\sum_{\alpha\in\mathbb{Z}^{2}}|b_{Q(\alpha)}|^{2}\leq\sum_{\alpha\in\mathbb{Z}^{2}}\frac{1}{|Q(\alpha)|}\int_{Q(\alpha)}|f|^{2}=\frac{1}{|Q(\alpha)|}\int_{\mathbb{R}^{2}}|f|^{2}.$
Here we use the fact that $|Q(\alpha)|=|Q|$ is independent of $\alpha$.
For the term $\sup_{x\in Q}|f(x)-b_{Q}|^{2}$, we write
$|f(x)-b_{Q}|=$ $\Big{|}f(x)-\frac{1}{|Q|}\int_{Q}f(y)\ \mathrm{d}y\Big{|}\leq\frac{1}{|Q|}\int_{Q}|f(x)-f(y)|\ \mathrm{d}y$
$=$ $\frac{1}{|Q|}\int_{I_{1}}\int_{I_{2}}|f(x_{1},x_{2})-f(y_{1},y_{2})|\ \mathrm{d}y_{2}\mathrm{d}y_{1}$
$\leq$ $\frac{1}{|Q|}\int_{I_{1}}\int_{I_{2}}|f(x_{1},x_{2})-f(y_{1},x_{2})|\ \mathrm{d}y_{2}\mathrm{d}y_{1}$
$+$ $\frac{1}{|Q|}\int_{I_{1}}\int_{I_{2}}|f(y_{1},x_{2})-f(y_{1},y_{2})|\ \mathrm{d}y_{2}\mathrm{d}y_{1}.$
Now we introduce the notations
[…]
Moreover, we have $x_{1},y_{1}\in I_{1}$ and
$|f(x_{1},x_{2})-f(y_{1},x_{2})|=\Big{|}\int_{x_{1}}^{y_{1}}\partial_{1}f(t,x_{2})\ \mathrm{d}t\Big{|}\leq\int_{I_{1}}|\partial_{1}f|(t,x_{2})\ \mathrm{d}t.$
The integrand $\partial_{1}f(t,x_{2})$ is independent of $y_{1}$, so by Cauchy–Schwarz,
[…]
For $Q(\alpha)$ with $\alpha\in\mathbb{Z}^{2}$, we write it as $I_{1}(\alpha_{1})\times I_{2}(\alpha_{2})$. Repeating the above argument and summing over $\alpha=(\alpha_{1},\alpha_{2})\in\mathbb{Z}^{2}$ shows that
$\sum_{\alpha\in\mathbb{Z}^{2}}\sup_{x\in Q(\alpha)}(I_{Q(\alpha)})^{2}\leq$ $\sum_{\alpha\in\mathbb{Z}^{2}}|I_{1}(\alpha_{1})|\int_{I_{1}(\alpha_{1})}\sup_{\begin{subarray}{c}x_{2}\in I_{2}(\alpha_{2})\end{subarray}}|\partial_{1}f|^{2}(t,x_{2})\,\mathrm{d}t$
$=$ $\sum_{\alpha_{1}\in\mathbb{Z}}|I_{1}(\alpha_{1})|\int_{I_{1}(\alpha_{1})}\sum_{\alpha_{2}\in\mathbb{Z}}\sup_{\begin{subarray}{c}x_{2}\in I_{2}(\alpha_{2})\end{subarray}}|\partial_{1}f|^{2}(t,x_{2})\,\mathrm{d}t$
$=$ $|I_{1}|\int_{\mathbb{R}}\sum_{\alpha_{2}\in\mathbb{Z}}\sup_{\begin{subarray}{c}x_{2}\in I_{2}(\alpha_{2})\end{subarray}}|\partial_{1}f|^{2}(t,x_{2})\,\mathrm{d}t.$
Here we use the fact that $|I_{1}(\alpha_{1})|=|I_{1}|$ is independent of $\alpha_{1}$, and the (disjoint) union $\bigcup_{\alpha_{1}\in\mathbb{Z}}I_{1}(\alpha_{1})$ is $\mathbb{R}$.
In [8], it was proved that for any interval $I$ and and smooth $f:I\rightarrow\mathbb{R}$,
$\sup_{I}|f|^{2}\leq 2\bigg{(}\frac{1}{|I|}\int_{I}|f|^{2}+|I|\int_{I}|f^{\prime}|^{2}\bigg{)}.$
Now apply this one-dimensional inequality to $I_{2}(\alpha_{2})$ and $x_{2}\mapsto\partial_{1}f(t,x_{2})$,
$\sup_{\begin{subarray}{c}x_{2}\in I_{2}(\alpha_{2})\end{subarray}}|\partial_{1}f|^{2}(t,x_{2})\leq 2\bigg{(}\frac{1}{|I_{2}(\alpha_{2})|}\int_{I_{2}(\alpha_{2})}|\partial_{1}f|^{2}(t,x_{2})\ \mathrm{d}x_{2}+|I_{2}(\alpha_{2})|\int_{I_{2}(\alpha_{2})}|\partial_{2}\partial_{1}f|^{2}(t,x_{2})\ \mathrm{d}x_{2}\bigg{)}.$
Summing over $\alpha_{2}$ gives
$\sum_{\alpha\in\mathbb{Z}^{2}}\sup_{x\in Q(\alpha)}(I_{Q(\alpha)})^{2}\leq\frac{2|I_{1}|}{|I_{2}|}\int_{\mathbb{R}}|\partial_{1}f|^{2}(t,x_{2})\ \mathrm{d}x_{2}+2|I_{1}||I_{2}|\int_{\mathbb{R}}|\partial_{2}\partial_{1}f|^{2}(t,x_{2})\ \mathrm{d}x_{2}.$
[…]
$|f(y_{1},x_{2})-f(y_{1},y_{2})|=\bigg{|}\int_{x_{2}}^{y_{2}}\partial_{2}f(y_{1},t)\ \mathrm{d}t\bigg{|}\leq\int_{I_{2}}|\partial_{2}f|(y_{1},t)\ \mathrm{d}t.$
[…]
$\frac{1}{|Q|}\int_{I_{1}}\int_{I_{2}}\int_{I_{2}}|\partial_{2}f|(y_{1},t)\ \mathrm{d}t\mathrm{d}y_{2}\mathrm{d}y_{1}$ $=\frac{1}{|I_{1}|}\iint_{I_{1}\times I_{2}}|\partial_{2}f|(y_{1},t)\,\mathrm{d}(y_{1},t)$
$\leq\frac{\sqrt{|Q|}}{|I_{1}|}\left(\int_{Q}|\partial_{2}f|^{2}\right)^{1/2},$
[…]
To summarize, we have proved that
$\sum_{\alpha}\sup_{Q(\alpha)}|f|^{2}$ $\lesssim\sum_{\alpha}\sup_{Q(\alpha)}|f-b_{Q(\alpha)}|^{2}+\sum_{\alpha}|b_{Q(\alpha)}|^{2}$
[…]
$\lesssim\frac{|I_{1}|}{|I_{2}|}\int|\partial_{1}f|^{2}+|Q|\int|\partial_{2}\partial_{1}f|^{2}+\frac{|I_{2}|}{|I_{1}|}\int|\partial_{2}f|^{2}+\frac{1}{|Q|}\int|f|^{2}.$
Now we back to (3.3) and suppose that $d=2$. An application of the Proposition 3.4 to the smooth function $f=\mathcal{F}(a\circ A^{k+N})$ and the rectangle $Q=R^{\ast}$ yields
$b^{2(k+N)}\sum_{\alpha\in\mathbb{Z}^{2}\setminus\{(0,0)\}}\sup_{\eta\in R^{\ast}(\alpha)}\Big{|}\mathcal{F}\big{(}a\circ A^{k+N}\big{)}(\eta)\Big{|}^{2}$
$\lesssim$ $b^{2(k+N)}\bigg{(}\frac{|I^{\ast}|}{|J^{\ast}|}\int|\partial_{1}f|^{2}+|R^{\ast}|\int|\partial_{2}\partial_{1}f|^{2}+\frac{|J^{\ast}|}{|I^{\ast}|}\int|\partial_{2}f|^{2}+\frac{1}{|R^{\ast}|}\int|f|^{2}\bigg{)}.$
By hypothesis, the atom $a$ is supported in $\Delta_{k}=A^{k}(\Delta_{0})$, so the function $a\circ A^{k+N}$ is supported in $A^{-N}(\Delta_{0})$. Using $\partial^{\alpha}\widehat{f}(\xi)=\mathcal{F}\big{(}(-2\pi\mathrm{i}x)^{\alpha}f(x)\big{)}(\xi)$ and the Plancherel’s identity,
$\frac{|I^{\ast}|}{|J^{\ast}|}\int|\partial_{1}\mathcal{F}(a\circ A^{k+N})|^{2}$ $\approx\frac{|I^{\ast}|}{|J^{\ast}|}\int_{A^{-N}(\Delta_{0})}|\eta_{1}|^{2}|a(A^{k+N}\eta)|^{2}\,\mathrm{d}\eta_{1}\mathrm{d}\eta_{2}$
$\lesssim\frac{|I^{\ast}|}{|J^{\ast}|\cdot|A^{k}(\Delta_{0})|^{2}}\int_{A^{-N}(\Delta_{0})}|\eta_{1}|^{2}\,\mathrm{d}\eta_{1}\mathrm{d}\eta_{2}$
$\lesssim_{N}\frac{|I^{\ast}|}{|J^{\ast}|b^{2k}}\int_{\Delta_{0}}|\eta_{1}|^{2}\,\mathrm{d}\eta_{1}\mathrm{d}\eta_{2}$
$\lesssim_{N,R^{\ast},\Delta_{0}}b^{-2k}.$
Therefore, the first term
$b^{2(k+N)}\frac{|I^{\ast}|}{|J^{\ast}|}\int|\partial_{1}f|^{2}\lesssim_{N,R^{\ast},\Delta_{0}}b^{2(k+N)}b^{-2k}\lesssim 1,$
where the implicit constant may depends on $b,N,\Delta_{0}$ and so on, but independent of $k$. Other terms can be treated by the same way. Similar argument gives the higher dimensional result of (3.3) for $d\geq 3$.
3.2. Proof of the necessary part
We are going to use the following well-known result on Bochner–Riesz multiplier.
Proposition 3.5.
Suppose that $\lambda>-1$. The function $m_{\lambda}\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ is defined by
$m_{\lambda}(\xi)\coloneqq\begin{cases}(1-|\xi|^{2})^{\lambda}&\quad\text{if }|\xi|\leq 1,\\ 0&\quad\text{if }|\xi|>1.\end{cases}$
The (inverse) Fourier transform of $m_{\lambda}$ can be written explicitly,
$m_{\lambda}^{\vee}(x)=\frac{\Gamma(\lambda+1)}{\pi^{\lambda}}\frac{J_{\frac{d}{2}+\lambda}(2\pi|x|)}{|x|^{\frac{d}{2}+\lambda}},$
where $J_{\frac{d}{2}+\lambda}$ is the first kind Bessel function of order $\frac{d}{2}+\lambda$. Moreover, for all $x\in\mathbb{R}^{d}$, the following inequality holds:
$m_{\lambda}^{\vee}(x)\leq\frac{C_{d,\lambda}}{1+|x|^{\frac{d+1}{2}+\lambda}},$
where $C_{d,\lambda}>0$ is a finite constant independent of $x$.
This result is standard and can be found in many textbook in Fourier analysis, for example, at the Appendix B of [5]. We remind the reader that $\Delta^{\ast}_{0}$ is the ellipse associated with the matrix $A^{\ast}$, so we may assume that $\Delta^{\ast}_{0}=P^{-1}(B(0,1))$ for certain $P\in\mathbf{GL}(d,\mathbb{R})$. Moreover, after rescaling, we can assume that $|\Delta_{0}^{\ast}|\approx 1$.
Lemma 3.6.
Let $g\in L^{2}(\mathbb{R}^{d})$ and assume that $\widehat{g}=0$ on $\Delta_{k}^{\ast}$. For $\lambda>-1$, we define
$\chi_{k,\lambda}(x)\coloneqq\begin{cases}(1-|P(A^{\ast})^{-k}x|^{2})^{\lambda}&\quad\text{if }|P(A^{\ast})^{-k}x|\leq 1,\\ 0&\quad\text{otherwise}.\end{cases}$
If $f\coloneq g\cdot\widehat{\chi_{k,\lambda}}$, then for $\lambda$ sufficiently large,
$\|f\|_{H^{1}}\lesssim_{\lambda,d}b^{k/2}\|g\|_{L^{2}}.$
Proof.
The argument here is adapted from [4]. We recall that $\Delta^{\ast}_{0}=\{x\in\mathbb{R}^{d}:|x|_{\ast}\leq 1\}=\{x\in\mathbb{R}^{d}:|Px|\leq 1\}$, hence $\mathrm{supp}(\chi_{k,\lambda})\subset\Delta^{\ast}_{k}$.
We begin by showing that $\int f=0$. This is easy. In fact, we have
$\widehat{f}(0)=\int_{\mathbb{R}^{d}}\widehat{g}(-y)\chi_{k,\lambda}(y)\,\mathrm{d}y=\int_{\Delta^{\ast}_{k}}\widehat{g}(-y)\chi_{k,\lambda}(y)\,\mathrm{d}y=0.$
To prove that $f\in H^{1}$, we use duality. Let $u\in\mathrm{BMO}$, then
$\Big{|}\int fu\Big{|}$ $=\Big{|}\int f(u-c)\Big{|}$ $\leq\|g\|_{2}\bigg{(}\int|u-c|^{2}|\widehat{\chi_{k,\lambda}}|^{2}\bigg{)}^{\frac{1}{2}},$
where $c$ is the average of $u$ over $A^{-k}\Delta_{0}$,
$c=\operatorname{Avg}_{A^{-k}(\Delta_{0})}\coloneqq\frac{1}{|A^{-k}\Delta_{0}|}\int_{A^{-k}\Delta_{0}}u(x)\,\mathrm{d}x.$
Easy calculation shows that
$\widehat{\chi_{k,\lambda}}(\xi)=\frac{b^{k}}{|\det(P)|}\widehat{m_{\lambda}}(P^{-\ast}A^{k}\xi).$
Using the properties of $m_{\lambda}$, we have
$|\widehat{\chi_{k,\lambda}}(\xi)|\lesssim\frac{b^{k}}{|\det(P)|}\times\frac{1}{1+|P^{-\ast}A^{k}\xi|^{\frac{d+1}{2}+\lambda}}.$
Since all norms on $\mathbb{R}^{d}$ are equivalent, we have $|P^{-\ast}A^{k}\xi|\approx_{P}|A^{k}\xi|$, and now
$\bigg{(}\int|u-c|^{2}|\widehat{\chi_{k,\lambda}}|^{2}\bigg{)}^{\frac{1}{2}}\lesssim_{P}b^{k}\bigg{(}\int\frac{|u-c|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}\,\mathrm{d}x\bigg{)}^{\frac{1}{2}}$
Decompose
$\int\frac{|u-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}$ $=\int_{A^{-k}(\Delta_{0})}\frac{|u-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}$
$+\sum_{l=0}^{\infty}\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}\frac{|u-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}$
For the integration on $A^{-k}(\Delta_{0})$, the trivial fact $|A^{k}x|\geq 0$ implies that
$\int_{A^{-k}(\Delta_{0})}\frac{|u-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}$ $\lesssim\int_{A^{-k}(\Delta_{0})}|u-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|^{2}$
$\lesssim|A^{-k}(\Delta_{0})|\cdot\|u\|_{\mathrm{BMO}}^{2}$
$\approx b^{-k}\cdot\|u\|_{\mathrm{BMO}}^{2}.$
We decompose, for $l\geq 0$, the integral $\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}$ as
$\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}\frac{|u-\operatorname{Avg}_{A^{l+1-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}$
$+$ $\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}\frac{|\operatorname{Avg}_{A^{l+1-k}(\Delta_{0})}(u)-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}.$
For $x\in A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})$, we have
$A^{k}(x)\in A^{l+1}(\Delta_{0})\setminus A^{l}(\Delta_{0}),\text{ and }\rho(A^{k}x)=b^{l}\geq 1.$
Proposition 2.1 says that $|x|\gtrsim\rho(x)^{\zeta_{-}}$ whenever $\rho(x)\geq 1$. Replacing $x$ by $A^{k}x$, we get $|A^{k}x|\gtrsim\rho(A^{k}x)^{\zeta_{-}}=b^{l\zeta_{-}}$.
Now
$\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}\frac{|u-\operatorname{Avg}_{A^{l+1-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}$
$\lesssim\frac{1}{b^{l(d+1+2\lambda)\zeta_{-}}}\int_{A^{l+1-k}(\Delta_{0})}|u-\operatorname{Avg}_{A^{l+1-k}(\Delta_{0})}(u)|^{2}$
$\lesssim\frac{|A^{l+1-k}(\Delta_{0})|}{b^{l(d+1+2\lambda)\zeta_{-}}}\|u\|_{\mathrm{BMO}}^{2}$
$\approx\frac{b^{l+1-k}}{b^{l(d+1+2\lambda)\zeta_{-}}}\|u\|_{\mathrm{BMO}}^{2}.$
For $u\in\mathrm{BMO}$, using the facts $\Delta_{0}\subset A(\Delta_{0})$ and $|A(\Delta_{0})|=b|\Delta_{0}|$,
$|\operatorname{Avg}_{\Delta_{0}}(u)-\operatorname{Avg}_{A(\Delta_{0})}(u)|$ $=\frac{1}{|\Delta_{0}|}\Big{|}\int_{\Delta_{0}}u-\operatorname{Avg}_{A(\Delta_{0})}(u)\Big{|}$
$\lesssim\frac{b}{|A(\Delta_{0})|}\int_{A(\Delta_{0})}\Big{|}u-\operatorname{Avg}_{A(\Delta_{0})}(u)\Big{|}$
$\leq b\|u\|_{\mathrm{BMO}}.$
Therefore, by telescoping,
$|\operatorname{Avg}_{A^{l+1-k}(\Delta_{0})}(u)-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|\lesssim(l+1)b\|u\|_{\mathrm{BMO}},$
and
$\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}\frac{|\operatorname{Avg}_{A^{l+1-k}(\Delta_{0})}(u)-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}$
$\lesssim$ $(l+1)^{2}b^{2}\|u\|_{\mathrm{BMO}}^{2}\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}\frac{1}{1+|A^{k}x|^{d+1+2\lambda}}$
$\lesssim$ $(l+1)^{2}b^{2}\|u\|_{\mathrm{BMO}}^{2}\frac{b^{l+1-k}}{b^{l(d+1+2\lambda)\zeta_{-}}}.$
Summing over $l$ gives
$\sum_{l=0}^{\infty}\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}\frac{|u-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}$
$\lesssim$ $\sum_{l=0}^{\infty}\frac{b^{l+1-k}(1+(l+1)^{2}b^{2})}{b^{l(d+1+2\lambda)\zeta_{-}}}\|u\|^{2}_{\mathrm{BMO}}.$
Now we choose $\lambda$ large, so that $1<(d+1+2\lambda)\zeta_{-}$, then the sum is controlled by $b^{-k}\|u\|^{2}_{\mathrm{BMO}}$. Combining these estimates gives us
$\Big{|}\int fu\Big{|}\lesssim b^{\frac{k}{2}}\|g\|_{2}\|u\|_{\mathrm{BMO}}.$
By duality, the $H^{1}$ norm of $f$ is controlled by $\mathcal{O}(b^{\frac{k}{2}}\|g\|_{2})$. ∎
Remark 3.
Under the isotropic settings, Sledd & Stegenga proved results in [8] by choosing the ball multiplier $m_{0}$. However, due to the anisotropic nature, the ball multiplier does not work here: We need an additional $\lambda>0$ to offset the increase of $b^{l}$ in the numerator. Readers can easily verify that, if $\lambda=0$, then the key inequality $1<(d+1+2\lambda)\zeta_{-}$ does not necessarily hold for a general dilation matrix $A$.
Now we back to the proof of the necessary part. Suppose that (3.1) holds. Applying (3.1) for $f=g\cdot\widehat{\chi_{k,\lambda}}$, and $\|g\|_{2}\leq 1$,
$\Big{|}\int\widehat{g}\ast\chi_{k,\lambda}\,\mathrm{d}\mu\Big{|}=\int|\widehat{f}|\,\mathrm{d}\mu\lesssim\|f\|_{H^{1}}\lesssim b^{\frac{k}{2}}.$
Notice that
$\int\left(\widehat{g}\ast\chi_{k,\lambda}\right)(x)\,\mathrm{d}\mu(x)=\int\widehat{g}(y)\int\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)\mathrm{d}y.$
Therefore
$\Big{|}\int\widehat{g}(y)\int\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)\mathrm{d}y\Big{|}\lesssim b^{\frac{k}{2}}.$
Recall that by our hypothesis, the Fourier transform $\widehat{g}$ is supported in $(\Delta^{\ast}_{k})^{c}$, so by duality,
$\int_{(\Delta^{\ast}_{k})^{c}}\bigg{|}\int\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)\bigg{|}^{2}\,\mathrm{d}y\lesssim b^{k}.$
Since $\mathrm{supp}(\chi_{k,\lambda})\subset\Delta^{\ast}_{k}$, we have $\int\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)=\int_{y+\Delta^{\ast}_{k}}\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)$. For $\epsilon>0$ small, and $x\in(1-\epsilon)\Delta^{\ast}_{k}$, it follows immediately that
$P(A^{\ast})^{-k}x\in(1-\epsilon)P(A^{\ast})^{-k}\Delta^{\ast}_{k}=(1-\epsilon)B(0,1),$
In consequence, the norm $|P(A^{\ast})^{-k}x|\leq 1-\epsilon$, and $\chi_{k,\lambda}(x)\gtrsim_{\epsilon,\lambda}1$.
Therefore, we have
$\int_{y+\Delta^{\ast}_{k}}\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)$ $\geq\int_{y+(1-\epsilon)\Delta^{\ast}_{k}}\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)$
$\gtrsim_{\epsilon,\lambda}\mu\big{(}y+(1-\epsilon)\Delta^{\ast}_{k}\big{)}$
This amounts to the following inequality:
$\int_{(\Delta^{\ast}_{k})^{c}}\mu\big{(}y+(1-\epsilon)\Delta^{\ast}_{k}\big{)}^{2}\,\mathrm{d}y\lesssim b^{k}.$
We know that $\Delta^{\ast}_{k}\subset(A^{\ast})^{k}(R^{\ast})$, so
$\bigcup_{\alpha\neq 0}(A^{\ast})^{k}(R^{\ast}(\alpha))=\Big{(}(A^{\ast})^{k}(R^{\ast})\Big{)}^{c}\subset(\Delta^{\ast}_{k})^{c},$
and
$\sum_{\alpha\neq 0}\int_{(A^{\ast})^{k}(R^{\ast}(\alpha))}\mu\big{(}y+(1-\epsilon)\Delta^{\ast}_{k}\big{)}^{2}\,\mathrm{d}y\lesssim b^{k}.$
We already knew that $(A^{\ast})^{-N}(R^{\ast})\subset\Delta^{\ast}_{0}$, so
$(1-\epsilon)(A^{\ast})^{-N}(A^{\ast})^{k}(R^{\ast})\subset(1-\epsilon)\Delta^{\ast}_{k}.$
By the geometry of $R^{\ast}$, it is easy to see that there exist a small constant $c>0$, so that
$c\cdot R^{\ast}\subset(1-\epsilon)(A^{\ast})^{-N}(R^{\ast}).$
Replacing $(1-\epsilon)\Delta^{\ast}_{k}$ by a smaller subset $(A^{\ast})^{k}(cR^{\ast})$ leads to
(3.4) $\sum_{\alpha\neq 0}\int_{(A^{\ast})^{k}(R^{\ast}(\alpha))}\mu\big{(}y+(A^{\ast})^{k}(cR^{\ast})\big{)}^{2}\,\mathrm{d}y\lesssim b^{k}.$
Next, we take a specific value $c=1/2$ as an example to explain the proof, the general case $c>0$ is similar, but the notations will be more complicated.
For $\alpha$ fixed, we decompose $R^{\ast}(\alpha)$ into 9 congruent parts $r_{i}$. It is evident that, the translated rectangle $y+\frac{1}{2}R^{\ast}$ contains $r_{i}$, whenever $y$ is near the center of $r_{i}$. Thus we may define a smaller rectangle concentric with $r_{i}$:
[…]
A trivial verification shows that $|r_{i}|/|r_{i}^{\prime}|\approx 1$. For general $c>0$, this area quotient depends of course on $c$, but the important point is that it is independent of $k$.
Figure 2. The rectangle $y+\frac{1}{2}R^{\ast}$ contains $r_{1}$, whenever $y\in r_{1}^{\prime}$.
[…]
Based on the previous analysis, the following inequality is clearly established:
$\int_{(A^{\ast})^{k}(R^{\ast}(\alpha))}\mu\big{(}y+(A^{\ast})^{k}(cR^{\ast})\big{)}^{2}\,\mathrm{d}y$ $=\sum_{1\leq i\leq 9}\int_{(A^{\ast})^{k}(r_{i})}\mu\big{(}y+(A^{\ast})^{k}(cR^{\ast})\big{)}^{2}\,\mathrm{d}y$
$\geq\sum_{1\leq i\leq 9}\int_{(A^{\ast})^{k}(r^{\prime}_{i})}\mu\big{(}y+(A^{\ast})^{k}(cR^{\ast})\big{)}^{2}\,\mathrm{d}y$
$\geq\sum_{1\leq i\leq 9}\int_{(A^{\ast})^{k}(r^{\prime}_{i})}\mu\big{(}(A^{\ast})^{k}(r_{i})\big{)}^{2}\,\mathrm{d}y$
$=\sum_{1\leq i\leq 9}\mu\big{(}(A^{\ast})^{k}(r_{i})\big{)}^{2}|(A^{\ast})^{k}(r^{\prime}_{i})|.$
We recall that $|r^{\prime}_{i}|\approx|r_{i}|=|R^{\ast}|/9$, and $|(A^{\ast})^{k}(r^{\prime}_{i})|=b^{k}|r^{\prime}_{i}|$. Thus the above inequality can be further written as
$\int_{(A^{\ast})^{k}(R^{\ast}(\alpha))}\mu\big{(}y+(A^{\ast})^{k}(cR^{\ast})\big{)}^{2}\,\mathrm{d}y$ $\gtrsim b^{k}|R^{\ast}|\sum_{1\leq i\leq 9}\mu\big{(}(A^{\ast})^{k}(r_{i})\big{)}^{2}$
$\gtrsim b^{k}|R^{\ast}|\mu\Big{(}\bigcup_{1\leq 1\leq 9}(A^{\ast})^{k}(r_{i})\Big{)}^{2}.$
The last line comes from the trivial estimate $\sum_{1\leq i\leq 9}|a_{i}|^{2}\geq\frac{1}{9}\big{(}\sum_{1\leq i\leq 9}|a_{i}|\big{)}^{2}$.
It is clear that
$\bigcup_{1\leq 1\leq 9}(A^{\ast})^{k}(r_{i})=(A^{\ast})^{k}\Big{(}\bigcup_{1\leq 1\leq 9}r_{i}\Big{)}=(A^{\ast})^{k}(R^{\ast}(\alpha)).$
Combining the above estimates gives us
$\int_{(A^{\ast})^{k}(R^{\ast}(\alpha))}\mu\big{(}y+(A^{\ast})^{k}(cR^{\ast})\big{)}^{2}\,\mathrm{d}y$ $\gtrsim b^{k}|R^{\ast}|\mu\bigg{(}(A^{\ast})^{k}(R^{\ast}(\alpha))\bigg{)}^{2}.$
Now back to (3.4), the following inequality holds:
$b^{k}$ $\gtrsim\sum_{\alpha\neq 0}\int_{(A^{\ast})^{k}(R^{\ast}(\alpha))}\mu\big{(}y+(A^{\ast})^{k}(cR^{\ast})\big{)}^{2}\,\mathrm{d}y$
$\gtrsim\sum_{\alpha\neq 0}b^{k}|R^{\ast}|\mu\bigg{(}(A^{\ast})^{k}(R^{\ast}(\alpha))\bigg{)}^{2}.$
This implies that, for all $k\in\mathbb{Z}$,
$\sum_{\alpha\neq 0}\mu\Big{(}(A^{\ast})^{k}\big{(}R^{\ast}(\alpha)\big{)}\Big{)}^{2}\lesssim|R^{\ast}|^{-1}\approx 1,$
which is exactly our desired result.
Remark 4.
In the original proof in [8], from the inequality
$\Big{|}\int\widehat{g}(y)\int\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)\mathrm{d}y\Big{|}\lesssim b^{\frac{k}{2}},$
the authors directly concluded that
$\int_{\mathbb{R}^{d}}\bigg{|}\int\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)\bigg{|}^{2}\,\mathrm{d}y\lesssim b^{k},$
and then (3.2) can be easily deduced.
Of course, this is inaccurate. Since $\widehat{g}=0$ in $\Delta_{k}^{\ast}$, we have no information on the behavior of $y\mapsto\int\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)$ for $y\in\Delta_{k}^{\ast}$, and we can only choose $\Omega=(\Delta_{k}^{\ast})^{c}$ (rather than $\mathbb{R}^{d}$) when using the duality:
$\|f\|_{L^{2}(\Omega)}=\sup_{\|g\|_{L^{2}(\Omega)}\leq 1}\left|\int fg\right|.$
4. Further results
Using the above method, few more results can be obtained. Under the isotropic settings, the following theorem was announced in [8].
The $H^{1}\mapsto L^{p}$ boundedness criterion.
Let $\mu$ be a positive Borel measure on $\mathbb{R}^{d}\setminus\{0\}$, and $H^{1}(\mathbb{R}^{d})$ be the isotropic Hardy space, then
$\left(\int_{\mathbb{R}^{d}}|\widehat{f}(\xi)|^{p}\,\mathrm{d}\mu(\xi)\right)^{\frac{1}{p}}\lesssim\|f\|_{H^{1}(\mathbb{R}^{d})},$
if and only if the following inequalities hold respectively:
  1. (1)
    When $1\leq p<2$,
    $\sup_{\epsilon>0}\Big{(}\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{0\}}\mu\big{(}Q_{\epsilon}(\alpha)\big{)}^{\frac{2}{2-p}}\Big{)}^{\frac{2-p}{2}}<\infty.$
  2. (2)
    When $2\leq p<\infty$,
    $\sup_{\epsilon>0}\mu\Big{(}\{x\in\mathbb{R}^{d}:\epsilon\leq|x|<2\epsilon\}\Big{)}<\infty.$
This theorem is stated in [8] but without proofs. We feel that, although the authors of [8] wrote that this was just an easy generalization of the $H^{1}$ boundedness criterion in Section 1, some modifications are still necessary and not that trivial. In this subsection, we state and give a sketchy proof of the corresponding result under the anisotropic settings.
Theorem 4.1.
Let $\mu$ be a positive Borel measure on $\mathbb{R}^{d}\setminus\{0\}$. Then
(4.1) $\left(\int_{\mathbb{R}^{d}}|\widehat{f}(\xi)|^{p}\,\mathrm{d}\mu(\xi)\right)^{\frac{1}{p}}\lesssim\|f\|_{H^{1}(\mathbb{R}^{d})},$
if and only if the following inequalities hold respectively.
  1. (1)
    When $1\leq p<2$,
    (4.2) $\sup_{k\in\mathbb{Z}}\left(\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{0\}}\mu\left(\left(A^{\ast}\right)^{k}\left(R^{\ast}\left(\alpha\right)\right)^{\frac{2}{2-p}}\right)\right)^{\frac{2-p}{2}}<\infty.$
  2. (2)
    When $2\leq p<\infty$,
    (4.3) $\sup_{k\in\mathbb{Z}}\mu\Big{(}\{x\in\mathbb{R}^{d}:\rho_{\ast}(x)=b^{k}\}\Big{)}<\infty.$
4.1. Proof of Theorem 4.1, when $1\leq p<2$
The proof at is somewhat similar to that of $p=1$.
The sufficient part
We first assume that (4.2) holds, it suffices to prove that for all $H^{1}$-atoms $a$,
$\int|\widehat{a}|^{p}\,\mathrm{d}\mu\lesssim 1.$
Same as before, we assume that $a$ is supported in $\Delta_{k}$, and we decompose the integral into two parts: integral over the region $(A^{\ast})^{-(k+N)}(R^{\ast})$, and over its complement.
Using Proposition 3.2 and Hölder’s inequality, we know that for $k\leq-(l+N)$,
$\mu\Big{(}(A^{\ast})^{l}(R^{\ast})\setminus(A^{\ast})^{l-N}(R^{\ast})\Big{)}$ $\leq\sum_{\alpha:\,\mathcal{O}_{N}(1)\textrm{ terms}}\mu\Big{(}(A^{\ast})^{l-N}(R^{\ast}(\alpha))\Big{)}$
$\lesssim\mathcal{O}_{N,p}(1)\bigg{(}\sum_{\alpha:\,\mathcal{O}_{N}(1)\textrm{ terms}}\mu\bigg{(}(A^{\ast})^{l-N}(R^{\ast}(\alpha))\bigg{)}^{\frac{2}{2-p}}\bigg{)}^{\frac{2-p}{2}}$
$\lesssim_{N,p}1.$
Now using Proposition 3.3, it is easy to prove that
$1$ $\gtrsim\int_{(A^{\ast})^{l}(R^{\ast})\setminus(A^{\ast})^{l-N}(R^{\ast})}\frac{1}{\rho_{\ast}^{p\zeta_{-}}(\xi)}\rho_{\ast}^{p\zeta_{-}}(\xi)\,\mathrm{d}\mu$
$\gtrsim\frac{1}{b^{(l+k+N)p\zeta_{-}}}\int_{(A^{\ast})^{l}(R^{\ast})\setminus(A^{\ast})^{l-N}(R^{\ast})}|\widehat{a}(\xi)|^{p}\,\mathrm{d}\mu.$
Summing over $\infty<l\leq-(k+N)$ yields that $\int_{(A^{\ast})^{-(k+N)}(R^{\ast})}|\widehat{a}|^{p}\,\mathrm{d}\mu\lesssim 1$.
For the integral over $\mathbb{R}^{d}\setminus(A^{\ast})^{-(k+N)}(R^{\ast})$, we just notice that
$\int_{\mathbb{R}^{d}\setminus(A^{\ast})^{-(k+N)}(R^{\ast})}|\widehat{a}|^{p}\,\mathrm{d}\mu$ $\leq\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{(0,0)\}}\mu\big{(}(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))\big{)}\sup_{(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))}|\widehat{a}|^{p}.$
Using (3.3), applying the Hölder’s inequality for the pair $(\frac{2}{2-p},\frac{2}{p})$ gives the desired result.
The necessary part
Now, on the contrary, we suppose that (4.1) holds. With a little effort, the following variant of Lemma 3.6 is not difficult to prove.
Lemma 4.2.
Suppose that $1<r<\infty$. Let $g\in L^{r}(\mathbb{R}^{d})$ and assume that $\widehat{g}=0$ on $\Delta_{k}^{\ast}$. If $f\coloneq g\cdot\widehat{\chi_{k,\lambda}}$, then for $\lambda$ sufficiently large,
$\|f\|_{H^{1}}\lesssim_{\lambda,d,r}b^{k/r}\|g\|_{L^{r}}.$
We omit the proof of this lemma since it is a step-by-step modification of that of Lemma 3.6. Regarding the meaning of “sufficiently large”, a careful calculation shows that any $\lambda$ strictly greater than $\frac{1}{r^{\prime}\zeta_{-}}-\frac{d+1}{2}$ works, where $r^{\prime}\coloneqq r/(r-1)$ is the conjugate exponent of $r$.
Now we apply $f=g\cdot\widehat{\chi_{k,\lambda}}$ to (4.1), it follows that $\left(\int|\widehat{g}\ast\chi_{k,\lambda}|^{p}\,\mathrm{d}\mu\right)^{\frac{1}{p}}\lesssim b^{\frac{k}{r}}\|g\|_{r}$. By duality, we know that for all $h\in L^{p^{\prime}}(\mathrm{d}\mu)$,
$\left|\int\left(\widehat{g}\ast\chi_{k,\lambda}\right)(x)\cdot h(x)\,\mathrm{d}\mu(x)\right|\lesssim b^{\frac{k}{r}}\|g\|_{r}\|h\|_{L^{p^{\prime}}(\mathrm{d}\mu)}.$
By Fubini’s theorem, it follows that
$\left|\int_{(\Delta^{\ast}_{k})^{c}}\widehat{g}(y)\int\chi_{k,\lambda}(x-y)h(x)\,\mathrm{d}\mu(x)\,\mathrm{d}y\right|\lesssim b^{\frac{k}{r}}\|g\|_{r}\|h\|_{L^{p^{\prime}}(\mathrm{d}\mu)}.$
Now we choose $r=2/(2-p)\geq 2$, since $1\leq p<2$. Hausdorff–Young inequality says that $\|g\|_{r}\leq\|\widehat{g}\|_{r^{\prime}}$, so
$\left|\int_{(\Delta^{\ast}_{k})^{c}}\widehat{g}(y)\int\chi_{k,\lambda}(x-y)h(x)\,\mathrm{d}\mu(x)\,\mathrm{d}y\right|\lesssim b^{\frac{k}{r}}\|\widehat{g}\|_{r^{\prime}}\|h\|_{L^{p^{\prime}}(\mathrm{d}\mu)}.$
By duality, we know that
$\left\|\int\chi_{k,\lambda}(x-y)h(x)\,\mathrm{d}\mu(x)\right\|_{L^{r}\left((\Delta^{\ast}_{k})^{c},\mathrm{d}y\right)}\lesssim b^{\frac{k}{r}}\|h\|_{L^{p^{\prime}}(\mathrm{d}\mu)},$
and again by duality,
$\left\|\left\|\chi_{k,\lambda}(x-y)\right\|_{L^{p}\left(\mathbb{R}^{d},\mathrm{d}\mu(x)\right)}\right\|_{L^{r}\left((\Delta^{\ast}_{k})^{c},\mathrm{d}y\right)}\lesssim b^{\frac{k}{r}}.$
Therefore, we have proved that
$\int_{(\Delta^{\ast}_{k})^{c}}\bigg{|}\int\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)\bigg{|}^{\frac{r}{p}}\,\mathrm{d}y\lesssim b^{k}.$
The rest part of the proof is similar to that of $p=1$.
4.2. Proof of Theorem 4.1, when $2\leq p<\infty$
This is somewhat different from the case $1\leq p<2$. When $d=1$ (hence isotropic) and $p=2$, Odysseas Bakas gives a proof in [1], Proposition 8, which works well under the anisotropic settings after minor modifications. The proof of this part is essentially due to Bakas.
We begin with a square function characterization of $H^{1}$. Let $\eta$ be a Schwartz function on $\mathbb{R}^{d}$ whose Fourier transform is nonnegative, supported in $(A^{\ast})^{2}(\Delta^{\ast})\setminus(A^{\ast})^{-1}(\Delta^{\ast})$, equal to $1$ on $(A^{\ast})(\Delta^{\ast})\setminus\Delta^{\ast}$. We define $\Psi$ via the Fourier transform:
$\widehat{\Psi}(\xi)\coloneqq\frac{\eta(\xi)}{\sum_{j\in\mathbb{Z}}\eta\left((A^{\ast})^{-j}\xi\right)}.$
It is easy to verify that $\widehat{\Psi}\in C^{\infty}_{c}(\mathbb{R}^{d})$, and $\widehat{\Psi}(\xi)\gtrsim 1$ on $(A^{\ast})(\Delta^{\ast})\setminus\Delta^{\ast}$, moreover,
$\sum_{j\in\mathbb{Z}}\widehat{\Psi}\left((A^{\ast})^{-j}\xi\right)=1,\text{ when }\xi\neq 0.$
Simple calculation shows that $\widehat{\Psi}\left((A^{\ast})^{-j}\xi\right)=\widehat{\Psi_{j}}(\xi)$, where $\Psi_{j}(x)\coloneqq b^{j}\Psi(A^{j}x)$. The following theorem is proved in [7], Theorem 2.8.
Theorem (Square function characterization of anisotropic $H^{1}$).
Assume that $f$ is a tempered distribution. The square function (or the g-function) of $f$ is defined by
$g(f)(x)\coloneqq\left(\sum_{j\in\mathbb{Z}}\left|\Psi_{j}\ast f\right|^{2}(x)\right)^{\frac{1}{2}}.$
If $f\in H^{1}$, then $\|f\|_{H^{1}}\approx\|g(f)\|_{L^{1}}$.
With these preparations in place, we can begin the proof.
The sufficient part
Assume that (4.2) holds, we decompose the integral into the sum
$\int_{\mathbb{R}^{d}}|\widehat{f}(\xi)|^{p}\,\mathrm{d}\mu(\xi)=\sum_{l\in\mathbb{Z}}\int_{(A^{\ast})^{l}(\Delta^{\ast})\setminus(A^{\ast})^{l-1}(\Delta^{\ast})}|\widehat{f}(\xi)|^{p}\,\mathrm{d}\mu(\xi).$
We fix an $l\in\mathbb{Z}$ and assume that $\xi\in(A^{\ast})^{l}(\Delta^{\ast})\setminus(A^{\ast})^{l-1}(\Delta^{\ast})$, which is equivalent to $\rho_{\ast}(\xi)=b^{l-1}$. It is not difficult to see that the sum $\sum_{j\in\mathbb{Z}}\widehat{\Psi}\left((A^{\ast})^{-j}\xi\right)$ is
$\sum_{l-2\leq j\leq l+2}\widehat{\Psi}\left((A^{\ast})^{-j}\xi\right),$
which has only 5 terms. So we can write $\widehat{f}(\xi)$ as
$\widehat{f}(\xi)\cdot 1=\widehat{f}(\xi)\sum_{l-2\leq j\leq l+2}\widehat{\Psi}\left((A^{\ast})^{-j}\xi\right)=\sum_{l-2\leq j\leq l+2}\widehat{\Psi_{j}\ast f}(\xi).$
Therefore,
$\left|\widehat{f}(\xi)\right|\leq\sum_{l-2\leq j\leq l+2}\left\|\widehat{\Psi_{j}\ast f}\right\|_{\infty}\leq\sum_{l-2\leq j\leq l+2}\left\|\Psi_{j}\ast f\right\|_{1}.$
Integrating over $(A^{\ast})^{l}(\Delta^{\ast})\setminus(A^{\ast})^{l-1}(\Delta^{\ast})$ shows that
$\int_{(A^{\ast})^{l}(\Delta^{\ast})\setminus(A^{\ast})^{l-1}(\Delta^{\ast})}|\widehat{f}|^{p}\,\mathrm{d}\mu$ $\leq\mu\left((A^{\ast})^{l}(\Delta^{\ast})\setminus(A^{\ast})^{l-1}(\Delta^{\ast})\right)\left(\sum_{l-2\leq j\leq l+2}\left\|\Psi_{j}\ast f\right\|_{1}\right)^{p}$
$\lesssim\sum_{l-2\leq j\leq l+2}\left\|\Psi_{j}\ast f\right\|_{1}^{p},$
where we used the hypothesis (4.3) and the trivial estimate $\left(\sum_{i=1}^{5}|a_{i}|\right)^{p}\lesssim\sum_{i=1}^{5}|a_{i}|^{p}$. Summing over $l\in\mathbb{Z}$ implies that
$\left(\int|\widehat{f}|^{p}\,\mathrm{d}\mu\right)^{\frac{1}{p}}\lesssim\left(\sum_{j\in\mathbb{Z}}\left\|\Psi_{j}\ast f\right\|_{1}^{p}\right)^{\frac{1}{p}}\lesssim\left(\sum_{j\in\mathbb{Z}}\left\|\Psi_{j}\ast f\right\|_{1}^{2}\right)^{\frac{1}{2}},$
where we used the condition $p\geq 2$ in the last inequality. Finally, the integral form of Minkowski’s inequality yields
$\left(\sum_{j\in\mathbb{Z}}\left\|\Psi_{j}\ast f\right\|_{1}^{2}\right)^{\frac{1}{2}}\leq\left\|\left(\sum_{j\in\mathbb{Z}}\left|\Psi_{j}\ast f\right|^{2}\right)^{\frac{1}{2}}\right\|_{1}\approx\|f\|_{H^{1}}.$
The necessary part
Assume that (4.1) holds. For $l\in\mathbb{Z}$, we choose $f\coloneqq\Psi_{-l}$. A direct calculation shows that
$\|f\|_{H^{1}}\approx\Big{\|}\Big{(}\sum_{-l-2\leq j\leq-l+2}|\Psi_{j}\ast f|^{2}\Big{)}^{\frac{1}{2}}\Big{\|}_{1}\approx\sum_{-l-2\leq j\leq-l+2}\|\Psi_{j}\ast f\|_{1}.$
After a change of variable, the last sum is
$\sum_{j=-2}^{2}\int\left|\int\Psi(z)\Psi(x-A^{j}z)\,\mathrm{d}z\right|\,\mathrm{d}x\approx 1.$
Therefore, the $H^{1}$ norm of $\Psi_{-l}$ is independent of $l$. However, from the construction of $\Psi$, we know that
$\widehat{\Psi_{-l}}(\xi)\gtrsim 1\text{ on }(A^{\ast})^{-l+1}(\Delta^{\ast})\setminus(A^{\ast})^{-l}(\Delta^{\ast}).$
So for $l\in\mathbb{Z}$,
$\mu\left((A^{\ast})^{-l+1}(\Delta^{\ast})\setminus(A^{\ast})^{-l}(\Delta^{\ast})\right)^{\frac{1}{p}}\lesssim\left(\int_{\mathbb{R}^{d}}|\widehat{\Psi_{-l}}(\xi)|^{p}\,\mathrm{d}\mu(\xi)\right)^{\frac{1}{p}}\lesssim\|\Psi_{-l}\|_{H^{1}(\mathbb{R}^{d})}\approx 1,$
where all implicit constants in $\lesssim$ are independent of $l$. Consequently, the inequality (4.3) holds.

### v3

@@
-An $H^{1}$ multiplier theorem on anisotropic Hardy sapces
+An $H^{1}$ multiplier theorem on anisotropic Hardy spaces
 Abstract.
-A parallel result of Sledd–Stegenga’s $H^{1}\rightarrow L^{1}$ multiplier theorem was obtained on the $H^{1}$ space under the anisotropic settings.
-Based on the same technique, an $H^{1}\rightarrow L^{p}$ multiplier theorem is also proved for $1\leq p<\infty$.
+We obtain a parallel result of Sledd–Stegenga’s $H^{1}\rightarrow L^{1}$ multiplier theorem under the anisotropic settings.
+Based on the same technique, we also prove an $H^{1}\rightarrow L^{p}$ multiplier theorem for $1\leq p<\infty$.
 1. Introduction
@@
 However, the anisotropic analog of Sledd–Stegenga’s criterion seems to be absent so far.
-We will close this gap.
+We close this gap in this note.
 The statement of the criterion under the anisotropic settings is more complicated and requires more preparation.
 Moreover, some arguments in [8] does not work any more, and we need few modifications.
-Also, there are problems in the original proofs of [8], which either omit many non-trivial steps or are too specific (only the 1-dimensional case is proved).
-Here we complete the missing details of the proof in [8].
+Also, there are gaps in the original proofs of [8], which either omit many non-trivial steps or are too specific (only the 1-dimensional case is proved, whereas the higher-dimensional case is not a direct result thereof).
+Here we complete the missing details in [8] incidentally.
+2. Acknowledgement
+This paper was completed at Nicolaus Copernicus University in Toruń, where the author was a postdoctoral researcher at that time.
+The author would like to thank Yuri Tomilov for introducing him to the paper [8].
+The author also express the gratitude to Marcing Bownik for introducing him the theory of anisotropic Hardy spaces.
 3. Preliminaries
-In this section, most of the text is adapted from the book [2].
+In this section, most of the text is adapted from [3].
 3.1. Dilation matrix and its homogeneous norm
-Let $A\in\mathbf{GL}(d,\mathbb{R})$ and $b\coloneqq|\det(A)|$. We say that $A$ is a dilation matrix, if all its eigenvalues $\{\lambda_{i}\}_{i=1}^{d}$ have modulus strictly greater than $1$.
+A matrix $A\in\mathbf{GL}(d,\mathbb{R})$ is a dilation matrix, if all its eigenvalues $\{\lambda_{i}\}_{i=1}^{d}$ have modulus strictly greater than $1$.
 For each dilation matrix $A$, there exists a measurable function (not unique) $\rho_{A}\colon\mathbb{R}^{d}\rightarrow[0,\infty)$, called a homogenous quasi-norm associated with $A$, satisfying the following properties:
@@
   2. (2)
-    $\rho_{A}(Ax)=b\rho_{A}(x)$ for all $x\in\mathbb{R}^{d}$.
+    $\rho_{A}(Ax)=b\rho_{A}(x)$ for all $x\in\mathbb{R}^{d}$, where $b\coloneqq|\det(A)|$.
   3. (3)
@@
 Although $\Delta$ is not unique, we will choose a specific one and refer it as the ellipse associated with $A$.
-The family of balls $\{\Delta_{k}\}_{k\in\mathbb{Z}}$ associated with $A$ around the origin is defined by
+The family of balls $\{\Delta_{k}\}_{k\in\mathbb{Z}}$ associated with AA around the origin is
 $\Delta_{k}\coloneqq A^{k}(\Delta),\text{ where }k\in\mathbb{Z}.$
 By Proposition 3.2, we have $\Delta_{k}\subset\Delta_{k+1}$, and $|\Delta_{k}|=b^{k}|\Delta_{0}|\approx b^{k}$.
-Although all quasi-norms $\rho_{A}$ are equivalent (see [2], Chapter 1, Lemma 2.4), we will use the following specific one: From the nested property of the family $\{\Delta_{k}\}_{k\in\mathbb{Z}}$, we know that for every $x\in\mathbb{R}^{d}\setminus\{0\}$, there exists a unique $k\in\mathbb{Z}$ so that $x\in\Delta_{k+1}\setminus\Delta_{k}$.
+Although all quasi-norms $\rho_{A}$ are equivalent (see [3], Chapter 1, Lemma 2.4), we will use the following specific one: From the nested property of the family $\{\Delta_{k}\}_{k\in\mathbb{Z}}$, for every $x\in\mathbb{R}^{d}\setminus\{0\}$, there exists unique $k\in\mathbb{Z}$ so that $x\in\Delta_{k+1}\setminus\Delta_{k}$.
 We define
@@
 The anisotropic dilation of $\varphi$ is defined by $\varphi_{k}(\cdot)\coloneqq b^{k}\varphi(A^{k}\cdot)$.
-For $f\in L^{1}(\mathbb{R}^{d})$, the smooth maximal function of $f$ with respect to $\varphi$ is defined by
+For $f\in L^{1}(\mathbb{R}^{d})$, the smooth maximal function of $f$ with respect to $\varphi$ is
 $\mathcal{M}_{\varphi}(f)(x)\coloneqq\sup_{k\in\mathbb{Z}}|f\ast\varphi_{k}|(x).$
-The anisotropic $H^{1}$ space associated with the dilation $A$ is defined by
+The anisotropic $H^{1}$ space associated with the dilation $A$ is
 $H^{1}_{A}(\mathbb{R}^{d})\coloneqq\Big\{f\in L^{1}(\mathbb{R}^{d}):\mathcal{M}_{\varphi}(f)\in L^{1}(\mathbb{R}^{d})\Big\}$
@@
     The function $a$ has vanishing moment: $\int a=0$.
-The atomic decomposition of $H^{1}_{A}$ states that $f\in H^{1}_{A}$ if and only if there exists atoms $\{a_{i}\}_{i=1}^{\infty}$ and a sequence $(\lambda_{i})_{i=1}^{\infty}\in\ell^{1}$, so that $f=\sum_{i=1}^{\infty}\lambda_{i}a_{i}$.
+The atomic decomposition of $H^{1}_{A}$ states that $f\in H^{1}_{A}$ if and only if there exist atoms $\{a_{i}\}_{i=1}^{\infty}$ and a sequence $(\lambda_{i})_{i=1}^{\infty}\in\ell^{1}$, so that $f=\sum_{i=1}^{\infty}\lambda_{i}a_{i}$.
 Moreover,
@@
 Theorem 4.1.
-Let $\mu$ be a positive Borel measure on $\mathbb{R}^{d}\setminus\{0\}$. Then
+Let $\mu$ be a positive Borel measure on $\mathbb{R}^{d}\setminus\{0\}$, then
 (4.1) $\int_{\mathbb{R}^{d}}|\widehat{f}|\,\mathrm{d}\mu\lesssim\|f\|_{H^{1}},$
@@
 Proposition 4.2.
-The region $(A^{\ast})^{l}(R^{\ast})\setminus(A^{\ast})^{l-N}(R^{\ast})$ can be covered by $\mathcal{O}_{N}(1)$ translations of $(A^{\ast})^{l-N}(R^{\ast}(\alpha))$ with $\alpha\neq(0,0)$.
+The region $(A^{\ast})^{l}(R^{\ast})\setminus(A^{\ast})^{l-N}(R^{\ast})$ can be covered by $\mathcal{O}_{N}(1)$ translations of $(A^{\ast})^{l-N}(R^{\ast}(\alpha))$ with $\alpha\neq 0$.
 This figure explains the proposition for $R^{\ast}\setminus(A^{\ast})^{-N}(R^{\ast})$.
@@
 $\lesssim_{N}1,$
-where we used the hypothesis (3.2) in the last line.
-We recall that $\Delta^{\ast}\subset R^{\ast}\subset(A^{\ast})^{N}(\Delta^{\ast})$. Applying the mapping $(A^{\ast})^{l-N}$ leads to
+where we used the hypothesis (4.2) in the last inequality.
+Since $\Delta^{\ast}\subset R^{\ast}\subset(A^{\ast})^{N}(\Delta^{\ast})$, applying the mapping $(A^{\ast})^{l-N}$ leads to
 $(A^{\ast})^{l-N}(R^{\ast})\subset(A^{\ast})^{l}(\Delta^{\ast})\subset(A^{\ast})^{l}(R^{\ast})\subset(A^{\ast})^{l+N}(\Delta^{\ast}),$
@@
 $|\widehat{a}(\xi)|\lesssim b^{k\zeta_{-}}\rho_{\ast}(\xi)^{\zeta_{-}},\text{ whenever }\rho_{\ast}(\xi)\leq b^{-k}.$
-Combining previous estimates, we can write
+Combining previous estimates,
 $1$ $\gtrsim\mu\Big((A^{\ast})^{l}(R^{\ast})\setminus(A^{\ast})^{l-N}(R^{\ast})\Big)$
@@
 $\lesssim 1.$
-The implicit constant depends (of course) on $N,b$ and $\zeta_{-}$, but not on $k$.
+The implicit constants are independent of $k$.
 Remark 1.
-In the original paper [8], the estimate over $(A^{\ast})^{-(k+N)}(R^{\ast})$ is omitted and the authors said it was an easy consequence of (3.2).
+In the original paper [8], the estimate over $(A^{\ast})^{-(k+N)}(R^{\ast})$ is omitted, and the authors said it was an easy consequence of (4.2).
 This is not quite accurate.
@@
 However, the correct condition is the sum over $\alpha\in\mathbb{Z}^{d}\setminus\{0\}$.
-In this case, we need few more work.
-Having completed the estimate of the integral on $(A^{\ast})^{-(k+N)}(R^{\ast})$, we now turn to control the integral on the complement $\mathbb{R}^{d}\setminus(A^{\ast})^{-(k+N)}(R^{\ast})$, it is easy to check that $\mathbb{R}^{d}\setminus R^{\ast}=\bigcup_{\alpha\in\mathbb{Z}^{d}\setminus\{(0,0)\}}R^{\ast}(\alpha)$, applying the invertible mapping $(A^{\ast})^{-(k+N)}$ gives
-$\mathbb{R}^{d}\setminus(A^{\ast})^{-(k+N)}(R^{\ast})=\bigcup_{\alpha\in\mathbb{Z}^{d}\setminus\{(0,0)\}}(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha)),$
-and the union is a disjoint union. Therefore
-$\int_{\mathbb{R}^{d}\setminus(A^{\ast})^{-(k+N)}(R^{\ast})}|\widehat{a}|\,\mathrm{d}\mu$ $=\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{(0,0)\}}\int_{(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))}|\widehat{a}|\,\mathrm{d}\mu$
+We now turn to control the integral on the complement $\mathbb{R}^{d}\setminus(A^{\ast})^{-(k+N)}(R^{\ast})$, it is easy to check that $\mathbb{R}^{d}\setminus R^{\ast}=\bigcup_{\alpha\in\mathbb{Z}^{d}\setminus\{0\}}R^{\ast}(\alpha)$, applying the invertible mapping $(A^{\ast})^{-(k+N)}$ gives
+$\mathbb{R}^{d}\setminus(A^{\ast})^{-(k+N)}(R^{\ast})=\bigcup_{\alpha\in\mathbb{Z}^{d}\setminus\{0\}}(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha)),$
+and the union is a disjoint union. Therefore,
+$\int_{\mathbb{R}^{d}\setminus(A^{\ast})^{-(k+N)}(R^{\ast})}|\widehat{a}|\,\mathrm{d}\mu$ $=\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{0\}}\int_{(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))}|\widehat{a}|\,\mathrm{d}\mu$
 […]
 By our hypothesis (4.2),
-$\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{(0,0)\}}\mu\Big{(}(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))\Big{)}^{2}\lesssim 1.$
+$\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{0\}}\mu\Big((A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))\Big)^{2}\lesssim 1.$
 By Cauchy–Schwarz, the proof will be completed if we can show that
-(3.3) $\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{(0,0)\}}\sup_{(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))}|\widehat{a}|^{2}\lesssim 1.$
+(4.3) $\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{0\}}\sup_{(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))}|\widehat{a}|^{2}\lesssim 1.$
 After a change of variable $\eta=(A^{\ast})^{k+N}\xi$, we have
@@
 We are reduced to control the sum
-$\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{(0,0)\}}\sup_{(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))}|\widehat{a}|^{2}=b^{2(k+N)}\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{(0,0)\}}\sup_{\eta\in R^{\ast}(\alpha)}\Big{|}\mathcal{F}\big{(}a\circ A^{k+N}\big{)}(\eta)\Big{|}^{2}$
+$\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{0\}}\sup_{(A^{\ast})^{-(k+N)}(R^{\ast}(\alpha))}|\widehat{a}|^{2}=b^{2(k+N)}\sum_{\alpha\in\mathbb{Z}^{d}\setminus\{0\}}\sup_{\eta\in R^{\ast}(\alpha)}\Big|\mathcal{F}\big(a\circ A^{k+N}\big)(\eta)\Big|^{2}$
 Proposition 4.4.
@@
 $\sum_{\alpha\in\mathbb{Z}}\sup_{I(\alpha)}|f|^{2}\leq 2\bigg(\frac{1}{|I|}\int_{\mathbb{R}}|f|^{2}+|I|\int_{\mathbb{R}}|f^{\prime}|^{2}\bigg).$
-In [8], the authors wrote that the higher-dimensional case is somewhat complicated so the proof was omitted.
+In [8], the authors wrote that the higher-dimensional case is somewhat complicated, so the proof was omitted.
 We give a proof of the two-dimensional case.
@@
 $\sup_{Q}|f|^{2}\leq 2\sup_{Q}|f-b_{Q}|^{2}+2|b_{Q}|^{2}.$
-Let us first see how to control $|b_{Q}|^{2}$, which is easier.
-The Cauchy–Schwarz inequality implies $|b_{Q}|^{2}\leq\frac{1}{|Q|}\int_{Q}|f|^{2}$.
-Summing over all $\alpha\in\mathbb{Z}^{2}$, we obtain
+Control $|b_{Q}|^{2}$: The Cauchy–Schwarz inequality implies $|b_{Q}|^{2}\leq\frac{1}{|Q|}\int_{Q}|f|^{2}$. Summing over all $\alpha\in\mathbb{Z}^{2}$, we obtain
 $\sum_{\alpha\in\mathbb{Z}^{2}}|b_{Q(\alpha)}|^{2}\leq\sum_{\alpha\in\mathbb{Z}^{2}}\frac{1}{|Q(\alpha)|}\int_{Q(\alpha)}|f|^{2}=\frac{1}{|Q(\alpha)|}\int_{\mathbb{R}^{2}}|f|^{2}.$
@@
 For the term $\sup_{x\in Q}|f(x)-b_{Q}|^{2}$, we write
-$|f(x)-b_{Q}|=$ $\Big{|}f(x)-\frac{1}{|Q|}\int_{Q}f(y)\ \mathrm{d}y\Big{|}\leq\frac{1}{|Q|}\int_{Q}|f(x)-f(y)|\ \mathrm{d}y$
-$=$ $\frac{1}{|Q|}\int_{I_{1}}\int_{I_{2}}|f(x_{1},x_{2})-f(y_{1},y_{2})|\ \mathrm{d}y_{2}\mathrm{d}y_{1}$
-$\leq$ $\frac{1}{|Q|}\int_{I_{1}}\int_{I_{2}}|f(x_{1},x_{2})-f(y_{1},x_{2})|\ \mathrm{d}y_{2}\mathrm{d}y_{1}$
-$+$ $\frac{1}{|Q|}\int_{I_{1}}\int_{I_{2}}|f(y_{1},x_{2})-f(y_{1},y_{2})|\ \mathrm{d}y_{2}\mathrm{d}y_{1}.$
-Now we introduce the notations
+[…]
+where we introduce the notations
 […]
@@
 […]
-Moreover, we have $x_{1},y_{1}\in I_{1}$ and
+Moreover, since $x_{1},y_{1}\in I_{1}$,
 $|f(x_{1},x_{2})-f(y_{1},x_{2})|=\Big|\int_{x_{1}}^{y_{1}}\partial_{1}f(t,x_{2})\ \mathrm{d}t\Big|\leq\int_{I_{1}}|\partial_{1}f|(t,x_{2})\ \mathrm{d}t.$
-The integrand $\partial_{1}f(t,x_{2})$ is independent of $y_{1}$, so by Cauchy–Schwarz,
+The integrand $\partial_{1}f(t,x_{2})$ is independent of $y_{1}$, by Cauchy–Schwarz,
 […]
@@
 Here we use the fact that $|I_{1}(\alpha_{1})|=|I_{1}|$ is independent of $\alpha_{1}$, and the (disjoint) union $\bigcup_{\alpha_{1}\in\mathbb{Z}}I_{1}(\alpha_{1})$ is $\mathbb{R}$.
-In [8], it was proved that for any interval $I$ and and smooth $f:I\rightarrow\mathbb{R}$,
+In [8], it was proved that for any interval $I$ and smooth $f:I\rightarrow\mathbb{R}$,
 $\sup_{I}|f|^{2}\leq 2\bigg(\frac{1}{|I|}\int_{I}|f|^{2}+|I|\int_{I}|f^{\prime}|^{2}\bigg).$
@@
 […]
-To summarize, we have proved that
+To summarize,
 $\sum_{\alpha}\sup_{Q(\alpha)}|f|^{2}$ $\lesssim\sum_{\alpha}\sup_{Q(\alpha)}|f-b_{Q(\alpha)}|^{2}+\sum_{\alpha}|b_{Q(\alpha)}|^{2}$
@@
 $\lesssim$ $b^{2(k+N)}\bigg(\frac{|I^{\ast}|}{|J^{\ast}|}\int|\partial_{1}f|^{2}+|R^{\ast}|\int|\partial_{2}\partial_{1}f|^{2}+\frac{|J^{\ast}|}{|I^{\ast}|}\int|\partial_{2}f|^{2}+\frac{1}{|R^{\ast}|}\int|f|^{2}\bigg).$
-By hypothesis, the atom $a$ is supported in $\Delta_{k}=A^{k}(\Delta_{0})$, so the function $a\circ A^{k+N}$ is supported in $A^{-N}(\Delta_{0})$.
+By hypothesis, the atom $a$ is supported in $\Delta_{k}=A^{k}(\Delta_{0})$, so the function $a\circ A^{k+N}$ is supported in $A^{-N}(\Delta_{0})$ and bounded by $|A^{-N}(\Delta_{0})|^{-1}$.
 Using $\partial^{\alpha}\widehat{f}(\xi)=\mathcal{F}\big((-2\pi\mathrm{i}x)^{\alpha}f(x)\big)(\xi)$ and the Plancherel’s identity,
@@
 $b^{2(k+N)}\frac{|I^{\ast}|}{|J^{\ast}|}\int|\partial_{1}f|^{2}\lesssim_{N,R^{\ast},\Delta_{0}}b^{2(k+N)}b^{-2k}\lesssim 1,$
-where the implicit constant may depends on $b,N,\Delta_{0}$ and so on, but independent of $k$.
+where implicit constants may depend on $b,N,\Delta_{0}$ and so on, but are independent of $k$.
 Other terms can be treated by the same way.
@@
 4.2. Proof of the necessary part
-We are going to use the following well-known result on Bochner–Riesz multiplier.
+We are going to use a well-known result on Bochner–Riesz multiplier. Suppose that $\lambda>-1$. The function $m_{\lambda}\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ is defined by
+$m_{\lambda}(\xi)\coloneqq\begin{cases}(1-|\xi|^{2})^{\lambda}&\quad\text{if }|\xi|\leq 1,\\ 0&\quad\text{if }|\xi|>1.\end{cases}$
 Proposition 4.5.
-Suppose that $\lambda>-1$. The function $m_{\lambda}\colon\mathbb{R}^{d}\rightarrow\mathbb{R}$ is defined by
-$m_{\lambda}(\xi)\coloneqq\begin{cases}(1-|\xi|^{2})^{\lambda}&\quad\text{if }|\xi|\leq 1,\\ 0&\quad\text{if }|\xi|>1.\end{cases}$
 The (inverse) Fourier transform of $m_{\lambda}$ can be written explicitly,
@@
 where $C_{d,\lambda}>0$ is a finite constant independent of $x$.
-This result is standard and can be found in many textbook in Fourier analysis, for example, at the Appendix B of [5].
-We remind the reader that $\Delta^{\ast}_{0}$ is the ellipse associated with the matrix $A^{\ast}$, so we may assume that $\Delta^{\ast}_{0}=P^{-1}(B(0,1))$ for certain $P\in\mathbf{GL}(d,\mathbb{R})$.
-Moreover, after rescaling, we can assume that $|\Delta_{0}^{\ast}|\approx 1$.
+This result is standard and can be found in many textbooks in Fourier analysis, for example, at the Appendix B of [5].
 Lemma 4.6.
@@
 The argument here is adapted from [4].
-We recall that $\Delta^{\ast}_{0}=\{x\in\mathbb{R}^{d}:|x|_{\ast}\leq 1\}=\{x\in\mathbb{R}^{d}:|Px|\leq 1\}$, hence $\mathrm{supp}(\chi_{k,\lambda})\subset\Delta^{\ast}_{k}$.
-We begin by showing that $\int f=0$. This is easy. In fact, we have
+We remind the reader that $\Delta^{\ast}_{0}$ is the ellipse associated with the matrix $A^{\ast}$, so we may assume that $\Delta^{\ast}_{0}=P^{-1}(B(0,1))$ for certain $P\in\mathbf{GL}(d,\mathbb{R})$.
+Moreover, after re-scaling, we can assume that $|\Delta_{0}^{\ast}|\approx 1$.
+It is trivial that $\mathrm{supp}(\chi_{k,\lambda})\subset\Delta^{\ast}_{k}$.
+We begin by showing that $\int f=0$. Using the support conditions of $\widehat{g}$ and $\chi_{k,\lambda}$,
 $\widehat{f}(0)=\int_{\mathbb{R}^{d}}\widehat{g}(-y)\chi_{k,\lambda}(y)\,\mathrm{d}y=\int_{\Delta^{\ast}_{k}}\widehat{g}(-y)\chi_{k,\lambda}(y)\,\mathrm{d}y=0.$
@@
 $\int\frac{|u-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}$ $=\int_{A^{-k}(\Delta_{0})}\frac{|u-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}$
-$+\sum_{l=0}^{\infty}\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}\frac{|u-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}$
-For the integration on $A^{-k}(\Delta_{0})$, the trivial fact $|A^{k}x|\geq 0$ implies that
+$+\sum_{l=0}^{\infty}\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}\frac{|u-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}.$
+The trivial fact $|A^{k}x|\geq 0$ implies that
 $\int_{A^{-k}(\Delta_{0})}\frac{|u-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}$ $\lesssim\int_{A^{-k}(\Delta_{0})}|u-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|^{2}$
@@
 $\approx b^{-k}\cdot\|u\|_{\mathrm{BMO}}^{2}.$
-We decompose, for $l\geq 0$, the integral $\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}$ as
+We decompose, for each $l\geq 0$, the integral $\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}$ as
 $\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}\frac{|u-\operatorname{Avg}_{A^{l+1-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}$
 $+$ $\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}\frac{|\operatorname{Avg}_{A^{l+1-k}(\Delta_{0})}(u)-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}.$
-For $x\in A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})$, we have
-$A^{k}(x)\in A^{l+1}(\Delta_{0})\setminus A^{l}(\Delta_{0}),\text{ and }\rho(A^{k}x)=b^{l}\geq 1.$
+For $x\in A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})$,
+$A^{k}x\in A^{l+1}(\Delta_{0})\setminus A^{l}(\Delta_{0}),\text{ and }\rho(A^{k}x)=b^{l}\geq 1.$
 Proposition 3.1 says that $|x|\gtrsim\rho(x)^{\zeta_{-}}$ whenever $\rho(x)\geq 1$. Replacing $x$ by $A^{k}x$, we get $|A^{k}x|\gtrsim\rho(A^{k}x)^{\zeta_{-}}=b^{l\zeta_{-}}$.
 Now
-$\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}\frac{|u-\operatorname{Avg}_{A^{l+1-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}$
-$\lesssim\frac{1}{b^{l(d+1+2\lambda)\zeta_{-}}}\int_{A^{l+1-k}(\Delta_{0})}|u-\operatorname{Avg}_{A^{l+1-k}(\Delta_{0})}(u)|^{2}$
+$\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}\frac{|u-\operatorname{Avg}_{A^{l+1-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}$ $\lesssim\int_{A^{l+1-k}(\Delta_{0})}\frac{|u-\operatorname{Avg}_{A^{l+1-k}(\Delta_{0})}(u)|^{2}}{b^{l(d+1+2\lambda)\zeta_{-}}}$
 $\lesssim\frac{|A^{l+1-k}(\Delta_{0})|}{b^{l(d+1+2\lambda)\zeta_{-}}}\|u\|_{\mathrm{BMO}}^{2}$
@@
 $\leq b\|u\|_{\mathrm{BMO}}.$
-Therefore, by telescoping,
-$|\operatorname{Avg}_{A^{l+1-k}(\Delta_{0})}(u)-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|\lesssim(l+1)b\|u\|_{\mathrm{BMO}},$
-and
+By telescoping,
+$|\operatorname{Avg}_{A^{l+1-k}(\Delta_{0})}(u)-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|\lesssim(l+1)b\|u\|_{\mathrm{BMO}}.$
+Consequently,
 $\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}\frac{|\operatorname{Avg}_{A^{l+1-k}(\Delta_{0})}(u)-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}$
@@
 $\lesssim$ $(l+1)^{2}b^{2}\|u\|_{\mathrm{BMO}}^{2}\frac{b^{l+1-k}}{b^{l(d+1+2\lambda)\zeta_{-}}}.$
-Summing over $l$ gives
+Summing over $l\in\mathbb{N}$,
 $\sum_{l=0}^{\infty}\int_{A^{l+1-k}(\Delta_{0})\setminus A^{l-k}(\Delta_{0})}\frac{|u-\operatorname{Avg}_{A^{-k}(\Delta_{0})}(u)|^{2}}{1+|A^{k}x|^{d+1+2\lambda}}$
@@
 Readers can easily verify that, if $\lambda=0$, then the key inequality $1<(d+1+2\lambda)\zeta_{-}$ does not necessarily hold for a general dilation matrix $A$.
-Now we back to the proof of the necessary part. Suppose that (3.1) holds. Applying (3.1) for $f=g\cdot\widehat{\chi_{k,\lambda}}$, and $\|g\|_{2}\leq 1$,
-$\Big{|}\int\widehat{g}\ast\chi_{k,\lambda}\,\mathrm{d}\mu\Big{|}=\int|\widehat{f}|\,\mathrm{d}\mu\lesssim\|f\|_{H^{1}}\lesssim b^{\frac{k}{2}}.$
-Notice that
-$\int\left(\widehat{g}\ast\chi_{k,\lambda}\right)(x)\,\mathrm{d}\mu(x)=\int\widehat{g}(y)\int\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)\mathrm{d}y.$
-Therefore
-$\Big{|}\int\widehat{g}(y)\int\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)\mathrm{d}y\Big{|}\lesssim b^{\frac{k}{2}}.$
-Recall that by our hypothesis, the Fourier transform $\widehat{g}$ is supported in $(\Delta^{\ast}_{k})^{c}$, so by duality,
+Back to the proof of the necessary part. Suppose that (4.1) holds. Applying (4.1) for $f=g\cdot\widehat{\chi_{k,\lambda}}$ with $\|g\|_{2}\leq 1$,
+$\Big|\int\widehat{g}\ast\chi_{k,\lambda}\,\mathrm{d}\mu\Big|=\int|\widehat{f}|\,\mathrm{d}\mu\lesssim\|f\|_{H^{1}}\lesssim b^{\frac{k}{2}},$
+while
+$\int\widehat{g}\ast\chi_{k,\lambda}\,\mathrm{d}\mu=\int\widehat{g}(y)\int\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)\mathrm{d}y.$
+Therefore,
+$\Big|\int\widehat{g}(y)\int\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)\mathrm{d}y\Big|\lesssim b^{\frac{k}{2}}$
+By our hypothesis, the Fourier transform $\widehat{g}$ is supported in $(\Delta^{\ast}_{k})^{c}$, so by duality,
 $\int_{(\Delta^{\ast}_{k})^{c}}\bigg|\int\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)\bigg|^{2}\,\mathrm{d}y\lesssim b^{k}.$
@@
 $P(A^{\ast})^{-k}x\in(1-\epsilon)P(A^{\ast})^{-k}\Delta^{\ast}_{k}=(1-\epsilon)B(0,1),$
-In consequence, the norm $|P(A^{\ast})^{-k}x|\leq 1-\epsilon$, and $\chi_{k,\lambda}(x)\gtrsim_{\epsilon,\lambda}1$.
-Therefore, we have
+In consequence, the norm $|P(A^{\ast})^{-k}x|\leq 1-\epsilon$, and $\chi_{k,\lambda}(x)\gtrsim_{\epsilon,\lambda}1$. Thus,
 $\int_{y+\Delta^{\ast}_{k}}\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)$ $\geq\int_{y+(1-\epsilon)\Delta^{\ast}_{k}}\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)$
-$\gtrsim_{\epsilon,\lambda}\mu\big{(}y+(1-\epsilon)\Delta^{\ast}_{k}\big{)}$
+$\gtrsim_{\epsilon,\lambda}\mu\big(y+(1-\epsilon)\Delta^{\ast}_{k}\big).$
 This amounts to the following inequality:
@@
 $\sum_{\alpha\neq 0}\int_{(A^{\ast})^{k}(R^{\ast}(\alpha))}\mu\big(y+(1-\epsilon)\Delta^{\ast}_{k}\big)^{2}\,\mathrm{d}y\lesssim b^{k}.$
-We already knew that $(A^{\ast})^{-N}(R^{\ast})\subset\Delta^{\ast}_{0}$, so
+Also, since $(A^{\ast})^{-N}(R^{\ast})\subset\Delta^{\ast}_{0}$,
 $(1-\epsilon)(A^{\ast})^{-N}(A^{\ast})^{k}(R^{\ast})\subset(1-\epsilon)\Delta^{\ast}_{k}.$
@@
 $c\cdot R^{\ast}\subset(1-\epsilon)(A^{\ast})^{-N}(R^{\ast}).$
-Replacing $(1-\epsilon)\Delta^{\ast}_{k}$ by a smaller subset $(A^{\ast})^{k}(cR^{\ast})$ leads to
+Replacing $(1-\epsilon)\Delta^{\ast}_{k}$ by the smaller subset $(A^{\ast})^{k}(cR^{\ast})$ leads to
 (4.4) $\sum_{\alpha\neq 0}\int_{(A^{\ast})^{k}(R^{\ast}(\alpha))}\mu\big(y+(A^{\ast})^{k}(cR^{\ast})\big)^{2}\,\mathrm{d}y\lesssim b^{k}.$
@@
 It is evident that, the translated rectangle $y+\frac{1}{2}R^{\ast}$ contains $r_{i}$, whenever $y$ is near the center of $r_{i}$.
-Thus we may define a smaller rectangle concentric with $r_{i}$:
+Thus, we define a smaller rectangle concentric with $r_{i}$:
 […]
-A trivial verification shows that $|r_{i}|/|r_{i}^{\prime}|\approx 1$. For general $c>0$, this area quotient depends of course on $c$, but the important point is that it is independent of $k$.
+A trivial verification shows that $|r_{i}|/|r_{i}^{\prime}|\approx 1$.
+For general $c>0$, the number of congruent parts into which $R^{\ast}(\alpha)$ is divided, and the area quotient $|r_{i}|/|r_{i}^{\prime}|$ depend of course on $c$, but the important point is that they are independent of $k$.
 Figure 2. The rectangle $y+\frac{1}{2}R^{\ast}$ contains $r_{1}$, whenever $y\in r_{1}^{\prime}$.
@@
 $=\sum_{1\leq i\leq 9}\mu\big((A^{\ast})^{k}(r_{i})\big)^{2}|(A^{\ast})^{k}(r^{\prime}_{i})|.$
-We recall that $|r^{\prime}_{i}|\approx|r_{i}|=|R^{\ast}|/9$, and $|(A^{\ast})^{k}(r^{\prime}_{i})|=b^{k}|r^{\prime}_{i}|$. Thus the above inequality can be further written as
+Since $|r^{\prime}_{i}|\approx|r_{i}|=|R^{\ast}|/9$, and $|(A^{\ast})^{k}(r^{\prime}_{i})|=b^{k}|r^{\prime}_{i}|$, the above inequality can be further written as
 $\int_{(A^{\ast})^{k}(R^{\ast}(\alpha))}\mu\big(y+(A^{\ast})^{k}(cR^{\ast})\big)^{2}\,\mathrm{d}y$ $\gtrsim b^{k}|R^{\ast}|\sum_{1\leq i\leq 9}\mu\big((A^{\ast})^{k}(r_{i})\big)^{2}$
@@
 $\int_{(A^{\ast})^{k}(R^{\ast}(\alpha))}\mu\big(y+(A^{\ast})^{k}(cR^{\ast})\big)^{2}\,\mathrm{d}y$ $\gtrsim b^{k}|R^{\ast}|\mu\bigg((A^{\ast})^{k}(R^{\ast}(\alpha))\bigg)^{2}.$
-Now back to (3.4), the following inequality holds:
+Now back to (4.4), the following inequality holds,
 $b^{k}$ $\gtrsim\sum_{\alpha\neq 0}\int_{(A^{\ast})^{k}(R^{\ast}(\alpha))}\mu\big(y+(A^{\ast})^{k}(cR^{\ast})\big)^{2}\,\mathrm{d}y$
@@
 5. Further results
-Using the above method, few more results can be obtained. Under the isotropic settings, the following theorem was announced in [8].
+Under the isotropic settings, the following theorem was announced in [8].
 The $H^{1}\mapsto L^{p}$ boundedness criterion.
@@
 Same as before, we assume that $a$ is supported in $\Delta_{k}$, and we decompose the integral into two parts: integral over the region $(A^{\ast})^{-(k+N)}(R^{\ast})$, and over its complement.
-Using Proposition 3.2 and Hölder’s inequality, we know that for $k\leq-(l+N)$,
+Using Proposition 4.2 and Hölder’s inequality, for $k\leq-(l+N)$,
 $\mu\Big((A^{\ast})^{l}(R^{\ast})\setminus(A^{\ast})^{l-N}(R^{\ast})\Big)$ $\leq\sum_{\alpha:\,\mathcal{O}_{N}(1)\textrm{ terms}}\mu\Big((A^{\ast})^{l-N}(R^{\ast}(\alpha))\Big)$
@@
 Regarding the meaning of “sufficiently large”, a careful calculation shows that any $\lambda$ strictly greater than $\frac{1}{r^{\prime}\zeta_{-}}-\frac{d+1}{2}$ works, where $r^{\prime}\coloneqq r/(r-1)$ is the conjugate exponent of $r$.
-Now we apply $f=g\cdot\widehat{\chi_{k,\lambda}}$ to (4.1), it follows that $\left(\int|\widehat{g}\ast\chi_{k,\lambda}|^{p}\,\mathrm{d}\mu\right)^{\frac{1}{p}}\lesssim b^{\frac{k}{r}}\|g\|_{r}$.
-By duality, we know that for all $h\in L^{p^{\prime}}(\mathrm{d}\mu)$,
+We apply $f=g\cdot\widehat{\chi_{k,\lambda}}$ to (5.1), it follows that $\left(\int|\widehat{g}\ast\chi_{k,\lambda}|^{p}\,\mathrm{d}\mu\right)^{\frac{1}{p}}\lesssim b^{\frac{k}{r}}\|g\|_{r}$.
+By duality, for all $h\in L^{p^{\prime}}(\mathrm{d}\mu)$,
 $\left|\int\left(\widehat{g}\ast\chi_{k,\lambda}\right)(x)\cdot h(x)\,\mathrm{d}\mu(x)\right|\lesssim b^{\frac{k}{r}}\|g\|_{r}\|h\|_{L^{p^{\prime}}(\mathrm{d}\mu)}.$
-By Fubini’s theorem, it follows that
+By Fubini’s theorem,
 $\left|\int_{(\Delta^{\ast}_{k})^{c}}\widehat{g}(y)\int\chi_{k,\lambda}(x-y)h(x)\,\mathrm{d}\mu(x)\,\mathrm{d}y\right|\lesssim b^{\frac{k}{r}}\|g\|_{r}\|h\|_{L^{p^{\prime}}(\mathrm{d}\mu)}.$
@@
 $\left|\int_{(\Delta^{\ast}_{k})^{c}}\widehat{g}(y)\int\chi_{k,\lambda}(x-y)h(x)\,\mathrm{d}\mu(x)\,\mathrm{d}y\right|\lesssim b^{\frac{k}{r}}\|\widehat{g}\|_{r^{\prime}}\|h\|_{L^{p^{\prime}}(\mathrm{d}\mu)}.$
-By duality, we know that
+By duality,
 $\left\|\int\chi_{k,\lambda}(x-y)h(x)\,\mathrm{d}\mu(x)\right\|_{L^{r}\left((\Delta^{\ast}_{k})^{c},\mathrm{d}y\right)}\lesssim b^{\frac{k}{r}}\|h\|_{L^{p^{\prime}}(\mathrm{d}\mu)},$
@@
 $\left\|\left\|\chi_{k,\lambda}(x-y)\right\|_{L^{p}\left(\mathbb{R}^{d},\mathrm{d}\mu(x)\right)}\right\|_{L^{r}\left((\Delta^{\ast}_{k})^{c},\mathrm{d}y\right)}\lesssim b^{\frac{k}{r}}.$
-Therefore, we have proved that
+Therefore,
 $\int_{(\Delta^{\ast}_{k})^{c}}\bigg|\int\chi_{k,\lambda}(x-y)\,\mathrm{d}\mu(x)\bigg|^{\frac{r}{p}}\,\mathrm{d}y\lesssim b^{k}.$
@@
 Let $\eta$ be a Schwartz function on $\mathbb{R}^{d}$ whose Fourier transform is nonnegative, supported in $(A^{\ast})^{2}(\Delta^{\ast})\setminus(A^{\ast})^{-1}(\Delta^{\ast})$, equal to $1$ on $(A^{\ast})(\Delta^{\ast})\setminus\Delta^{\ast}$.
-We define $\Psi$ via the Fourier transform:
+We define $\Psi$ via the Fourier transform,
 $\widehat{\Psi}(\xi)\coloneqq\frac{\eta(\xi)}{\sum_{j\in\mathbb{Z}}\eta\left((A^{\ast})^{-j}\xi\right)}.$
@@
 $\widehat{f}(\xi)\cdot 1=\widehat{f}(\xi)\sum_{l-2\leq j\leq l+2}\widehat{\Psi}\left((A^{\ast})^{-j}\xi\right)=\sum_{l-2\leq j\leq l+2}\widehat{\Psi_{j}\ast f}(\xi).$
-Therefore,
-$\left|\widehat{f}(\xi)\right|\leq\sum_{l-2\leq j\leq l+2}\left\|\widehat{\Psi_{j}\ast f}\right\|_{\infty}\leq\sum_{l-2\leq j\leq l+2}\left\|\Psi_{j}\ast f\right\|_{1}.$
+Taking supremum over $\xi$,
+$\sup_{\xi\in(A^{\ast})^{l}(\Delta^{\ast})\setminus(A^{\ast})^{l-1}(\Delta^{\ast})}\left|\widehat{f}(\xi)\right|\leq\sum_{l-2\leq j\leq l+2}\left\|\widehat{\Psi_{j}\ast f}\right\|_{\infty}\leq\sum_{l-2\leq j\leq l+2}\left\|\Psi_{j}\ast f\right\|_{1}.$
 Integrating over $(A^{\ast})^{l}(\Delta^{\ast})\setminus(A^{\ast})^{l-1}(\Delta^{\ast})$ shows that
 $\int_{(A^{\ast})^{l}(\Delta^{\ast})\setminus(A^{\ast})^{l-1}(\Delta^{\ast})}|\widehat{f}|^{p}\,\mathrm{d}\mu$ $\leq\mu\left((A^{\ast})^{l}(\Delta^{\ast})\setminus(A^{\ast})^{l-1}(\Delta^{\ast})\right)\left(\sum_{l-2\leq j\leq l+2}\left\|\Psi_{j}\ast f\right\|_{1}\right)^{p}$
-$\lesssim\sum_{l-2\leq j\leq l+2}\left\|\Psi_{j}\ast f\right\|_{1}^{p},$
-where we used the hypothesis (4.3) and the trivial estimate $\left(\sum_{i=1}^{5}|a_{i}|\right)^{p}\lesssim\sum_{i=1}^{5}|a_{i}|^{p}$. Summing over $l\in\mathbb{Z}$ implies that
-$\left(\int|\widehat{f}|^{p}\,\mathrm{d}\mu\right)^{\frac{1}{p}}\lesssim\left(\sum_{j\in\mathbb{Z}}\left\|\Psi_{j}\ast f\right\|_{1}^{p}\right)^{\frac{1}{p}}\lesssim\left(\sum_{j\in\mathbb{Z}}\left\|\Psi_{j}\ast f\right\|_{1}^{2}\right)^{\frac{1}{2}},$
-where we used the condition $p\geq 2$ in the last inequality. Finally, the integral form of Minkowski’s inequality yields
+$\lesssim\sum_{l-2\leq j\leq l+2}\left\|\Psi_{j}\ast f\right\|_{1}^{p}.$
+where we used the hypothesis (5.3) and the trivial estimate $\left(\sum_{i=1}^{5}|a_{i}|\right)^{p}\lesssim\sum_{i=1}^{5}|a_{i}|^{p}$. Summing over $l\in\mathbb{Z}$, and use the condition $p\geq 2$,
+$\left(\int|\widehat{f}|^{p}\,\mathrm{d}\mu\right)^{\frac{1}{p}}\lesssim\left(\sum_{j\in\mathbb{Z}}\left\|\Psi_{j}\ast f\right\|_{1}^{p}\right)^{\frac{1}{p}}\lesssim\left(\sum_{j\in\mathbb{Z}}\left\|\Psi_{j}\ast f\right\|_{1}^{2}\right)^{\frac{1}{2}}.$
+Finally, the integral form of Minkowski’s inequality yields
 $\left(\sum_{j\in\mathbb{Z}}\left\|\Psi_{j}\ast f\right\|_{1}^{2}\right)^{\frac{1}{2}}\leq\left\|\left(\sum_{j\in\mathbb{Z}}\left|\Psi_{j}\ast f\right|^{2}\right)^{\frac{1}{2}}\right\|_{1}\approx\|f\|_{H^{1}}.$
@@
 $\sum_{j=-2}^{2}\int\left|\int\Psi(z)\Psi(x-A^{j}z)\,\mathrm{d}z\right|\,\mathrm{d}x\approx 1.$
-Therefore, the $H^{1}$ norm of $\Psi_{-l}$ is independent of $l$. However, from the construction of $\Psi$, we know that
+Therefore, the $H^{1}$ norm of $\Psi_{-l}$ is independent of $l$. From the construction of $\Psi$,
 $\widehat{\Psi_{-l}}(\xi)\gtrsim 1\text{ on }(A^{\ast})^{-l+1}(\Delta^{\ast})\setminus(A^{\ast})^{-l}(\Delta^{\ast}).$